### v2

# Automating Governing Knowledge Commons and Contextual Integrity (GKC-CI) Privacy Policy Annotations with Large Language Models

Abstract.
Identifying contextual integrity (CI) and governing knowledge commons (GKC) parameters in privacy policy texts can facilitate normative privacy analysis. However, GKC-CI annotation has heretofore required manual or crowdsourced effort. This paper demonstrates that high-accuracy GKC-CI parameter annotation of privacy policies can be performed automatically using large language models. We fine-tune 50 open-source and proprietary models on 21,588 GKC-CI annotations from 16 ground truth privacy policies. Our best performing model has an accuracy of 90.65%, which is comparable to the accuracy of experts on the same task. We apply our best performing model to 456 privacy policies from a variety of online services, demonstrating the effectiveness of scaling GKC-CI annotation for privacy policy exploration and analysis. We publicly release our model training code, training and testing data, an annotation visualizer, and all annotated policies for future GKC-CI research.
privacy, contextual integrity, governing knowledge commons, natural language processing, large language model, text tagging
1. Introduction
Privacy policies are notoriously complex and lengthy documents (McDonald and Cranor, 2008). These policies are often written in complex language or “legalese” to obfuscate the extent of data collection and discourage consumers from closely interrogating their privacy implications (Jensen and Potts, 2004; Reidenberg et al., 2016; Amos et al., 2021). Most consumers therefore choose to ignore privacy policies when agreeing to online terms and services (Rudolph et al., 2018). Even experts have difficulty interpreting some privacy policies (Reidenberg et al., 2015). However, privacy policies remain essential to Internet privacy broadly and to the privacy-relevant behaviors of online services.
The continued importance of privacy policies has motivated substantial research into structured methods of privacy policy analysis. Some of these methods seek to provide clearer or more easily digestible information to consumers or developers (Sadeh et al., 2013; Angulo et al., 2012; Brodie et al., 2006; Wilson et al., 2016a), while others facilitate academic studies of the policies themselves, their relation to company behavior, or to privacy regulation (Liu et al., 2014; Winkler and Zeadally, 2016; Andow et al., 2019; Perez et al., 2018; Andow et al., 2020; Tesfay et al., 2018). Both approaches often employ annotation—labeling relevant parts of privacy policy texts with metadata—as a primary technique.
Early successful efforts involved annotating privacy policies with a large set of metadata tags (Wilson et al., 2016a). A more recent approach (Shvartzshnaider et al., 2019a) has leveraged the theory of contextual integrity (CI) (Nissenbaum, 2009) to annotate privacy policies. CI annotation uses a small set of theoretically grounded tags to facilitate comparative and longitudinal analysis of data handling practices and policy ambiguities (Shvartzshnaider et al., 2019a). Research using CI as a theoretical foundation is widespread across topics in security and privacy, human-computer interaction, formal modeling, and legal analysis (Brehm and Shvartzhnaider, 2023; Apthorpe et al., 2019; Benthall, 2022; Shvartzshnaider et al., 2019b; Wijesekera et al., 2015; Datta et al., 2011). Unlike differential privacy and other probabilistic privacy definitions, CI extends a normative notion of privacy that is a more natural fit for natural language documents like privacy policies (Nissenbaum, 2009).
Recent work has shown that CI is even more effective if expanded using the governing knowledge commons framework (GKC) (Sanfilippo et al., 2018; Frischmann et al., 2014). GKC provides an institutional grammar for describing strategies, norms, and rules around shared knowledge resources. The unified GKC-CI framework (Shvartzshnaider et al., 2022) (Section 3) enables straightforward identification of privacy policy ambiguities that reduce interpretability and provide excessive leeway for behavior users may consider privacy-violating. GKC-CI also enables normative analyses of contextual information transfers, the rules-in-use, and the rules-on-the-books that govern data handling practices.
All previous uses of CI parameter annotation for privacy policy analysis have involved human effort by experts or crowdworkers. Manual annotation by expert researchers produces high-quality results, but the process is tedious and slow. Crowdsourcing produces annotations more quickly, but there is a significant rate of poor-quality annotations since the annotation task is inherently nuanced  (Shvartzshnaider et al., 2019a). Combining multiple crowdsourced annotations through a voting process can improve overall performance but further increases expense, as multiple crowdworkers must be hired to annotate overlapping sections of privacy policy text (Shvartzshnaider et al., 2019a). A prior study spent approximately \$200 for crowdworker annotation of only 48 excerpts from 16 privacy policies (Shvartzshnaider et al., 2019a).
Furthermore, no previous work on automating privacy policy annotation with machine learning (Wilson et al., 2018; Harkous et al., 2018; Bannihatti Kumar et al., 2020; Andow et al., 2020, 2019; Zimmeck et al., n. d.; Story et al., 2019; Tang et al., 2023) has been based on CI or GKC-CI. Rather, this prior work has focused on different sets of annotation tags that are not useful to the GKC/CI scholarly community as they do not produce results relevant to GKC/CI analysis. This motivates the development of novel automated techniques for the GKC-CI annotation task.
In this paper, we train a variety of large language models (LLMs) to perform automated GKC-CI parameter annotations of privacy policies. Doing so demonstrates the feasibility of large-scale longitudinal and cross-industry analysis of privacy policies using the GKC-CI framework. Specifically, we train and evaluate 50 LLMs from five different model families, ranging from open-source to proprietary models. We perform two ablation studies with the LLMs, and briefly remark on some key lessons which we believe may be broadly applicable to other researchers working in this space. We then proceed to a nuanced discussion of the behavior of our best performing model, particularly with respect to the model’s errors.
We observe that of the 50 models we benchmark, a version of GPT-3.5 Turbo performs the best. We find that the model boasts a robust accuracy of 90.65%, better than that of prior crowdsourcing approaches (Shvartzshnaider et al., 2019a). We consider an accurate annotation to be an exact string match of the human annotator’s text. In contrast to prior approaches, our best performing LLM only costs \$0.23-0.44 on average to annotate a privacy policy (depending on length) and is able to annotate an entire privacy policy in one minute or less on average (~5000 words a minute). We could annotate all policy excerpts from (Shvartzshnaider et al., 2019a) for only \$7 and in less than one minute.
We use our best performing model to annotate longitudinal and cross-industry policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). We evaluate these policies based on GKC-CI parameter counts, densities, and variances, demonstrating how GKC-CI annotation can be used to highlight policies of interest for further analysis. We also present a Python-based GUI that visualizes the results of GKC-CI annotation, showing privacy policy text with annotated parameters highlighted by color.
In summary, this paper presents a practical method for analyzing privacy policies that integrates machine learning, privacy law, and governance. This is the first automated method for GKC-CI parameter annotation in legal documents, which will be extremely useful to the growing GKC and CI research communities. This approach could also encourage collaboration and open up new research opportunities in these interconnected fields, promoting more in-depth exploration of privacy protection and compliance. More specifically,
  • We demonstrate that accurate GKC-CI parameter annotations of privacy policies can be performed automatically by a fine-tuned Large-Language Model (LLM), substantially improving scalability and reducing expense compared to manual and crowdsourcing approaches. Further, through training data review, we find these annotations are as accurate as those obtained from humans.
  • We train and publicly release the data and scripts necessary to reproduce a large language model capable of performing automatic GKC-CI annotation of privacy policies. Our automated approach has an average per-policy annotation cost between \$0.23-0.44, depending on the length of the policy. [^1] The cost estimation range is the average cost of annotating privacy policies from both the 164 least popular and the 164 most popular websites, per their Tranco (Pochat et al., 2018) ranking, in the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). We note that reproducing our model currently incurs a one-time training cost of \$195.
  • We perform a large-scale longitudinal and cross-industry analysis of privacy policies using our best performing model. We demonstrate that the annotations can highlight policies with atypical parameter densities and distributions that may be good candidates for future in-depth evaluation. We compile all 456 annotated policies into a GitHub Repository, which we make publicly available. [^2] https://github.com/JakeC007/Automated_GKC-CI_Privacy_Policy_Annotations
2. Related Work
Substantial prior research has focused on systematic analyses of privacy policies. These analyses were often intended to improve consumer understanding of data handling processes and facilitate academic study of Internet privacy trends. This paper builds on this foundation, contributing to the broad goal of developing a library of effective, scalable, and inexpensive privacy policy analysis techniques suitable for a range of applications.
2.1. The Usable Privacy Project
The Usable Privacy Project (Sadeh et al., 2013) from Carnegie Mellon University is perhaps the most visibly successful application of annotation as a method for privacy policy interpretation and explanation. This project started in 2016 with a study by Wilson et al. (Wilson et al., 2016a), that recruited law students to manually annotate privacy policies with metadata tags such as “first party collection/use,” “user choice/control,” “data retention,” and “data security.” Wilson et al. also showed (Wilson et al., 2016b) that annotations produced by crowdworkers agreed with those of expert annotators over 80% of the time. This showed that crowdsourcing techniques could be used to identify paragraphs describing specific data handling practices in privacy policies.
In 2018, Wilson et al. used 115 expert-labeled policies to train logistic regression, support vector machine, and convolutional neural network models to automatically label sentences or segments of privacy policies with data practice categories (Wilson et al., 2018). Their best models had average F1 scores of $0.66$ for policy sentences and $0.78$ for policy segments. These techniques have been applied to over 7000 privacy policies from 2017, with results posted on the Usable Privacy Project website to inform consumers of the wide variety of information handling practices conducted by online services. [^3] https://explore.usableprivacy.org/
This line of research, while independently impressive, is not relevant for the community of scholars using contextual integrity or governing knowledge commons to interpret privacy policies and other legal documents. The sets of annotation tags used for the Usable Privacy Project are not analogous to the 5-parameter information flow description fundamental to CI nor the institutional grammar fundamental to GKC. This means that a separate thread of research, described below, has been needed to develop privacy policy annotation approaches useful for GKC/CI analyses.
2.2. Manual CI Annotation
Unlike the previous work by Wilson et al. (Wilson et al., 2016a, b, 2018) and the Usable Privacy Project, our work is based on Shvartzshnaider et al.’s GKC-CI framework (Shvartzshnaider et al., 2022), which provides a theoretically grounded basis for identifying ambiguity and potential privacy-violating behavior. Our work is less focused on helping consumers understand privacy policies than (Wilson et al., 2016a, b, 2018) and is more focused on automatable, longitudinal, and cross-industry analysis of privacy policies. Our work is also novel in its use of large-language models (the current state-of-the-art in natural language processing) and achieves considerably better performance than the machine learning annotation in (Wilson et al., 2018).
In 2019, Shvartzshnaider et al. (Shvartzshnaider et al., 2019a) used the theory of contextual integrity (CI) (Nissenbaum, 2009) to inform a new approach to privacy policy annotation. This approach seeks to identify the five information flow parameters defined by CI (Section 3) in privacy policy text. CI parameter annotation enabled the identification of ambiguities in information transfer descriptions. Shvartzshnaider et al. (Shvartzshnaider et al., 2022) combined contextual integrity with governing knowledge commons (GKC) (Frischmann et al., 2014; Sanfilippo et al., 2018) in 2022 to create a combined GKC-CI framework. GKC-CI extends the potential scope of CI annotation to eight total parameters, four from CI and four from the GKC institutional grammar (Section 3).
While Shvartzshnaider et al. (Shvartzshnaider et al., 2019a) successfully motivated CI parameter annotation for privacy policy analysis, questions of scalability remained. As with most annotation tasks, manual annotation by experts is highly accurate but tedious and slow. Shvartzshnaider et al. demonstrated that crowdsourcing could partially solve this problem but remains expensive, as high error rates necessitated the combination of multiple overlapping crowdsourced annotations per policy segment to increase precision. The resulting crowdsourced annotations still had a relatively high rate of false negative errors, i.e., parameters missed by the majority of crowdworkers.
The scalability issues posed by crowdworker annotation clearly motivate this study, which seeks to automate CI parameter annotation through the use of large language models (LLMs). Our work is also novel in its use of the expanded eight-parameter GKC-CI labels as annotation tags (Section 3) rather than the five-parameter CI tags used in Shvartzshnaider et al.’s original paper.
We apply our automated method to annotate a large corpus of privacy policies, including up to 20 years of longitudinal policies from 8 major technology companies (128 policies) and 328 contemporary policies from across the technology industry. This is orders of magnitude more privacy policies than have been manually annotated for GKC-CI research in previous work (Shvartzshnaider et al., 2019a).
2.3. Privacy Policy Analysis With Machine Learning
Several other studies have also applied machine learning to privacy policies, although none have been based on nor produced output useful for GKC-CI research. In 2018, Harkous et al. (Harkous et al., 2018) trained a hierarchy of convolutional neural networks to build a Question-Answering system that supports free-form querying of privacy policy content. Other ML-based approaches essentially parse privacy policies for information of interest, such as by using a logistic regression model to identify opt-out statements in privacy policy text (Bannihatti Kumar et al., 2020). PoliCheck (Andow et al., 2020), an expansion of PolicyLint (Andow et al., 2019), is capable of differentiating between first-party and third-party entities in flow-to-policy consistency analysis. Zimmeck et al. (Zimmeck et al., n. d.) and Story et al. (Story et al., 2019) used support vector machines to identify non-compliance between Android application code and the applications’ privacy policies. Their approach could be used to highlight these statements for consumers to make opt-out decisions without needing to read the entire policy themselves. Our application of machine learning to GKC-CI privacy policy annotation is similarly tightly focused, but on a task that does not overlap these earlier works.
More recently, some research groups have utilized LLMs broadly in the legal space or more specifically for privacy policy analysis. LLM benchmarks in the legal space include that of (Dai et al., 2023) and (Fei et al., 2023). Further, Ravichander et al. (Ravichander et al., 2019) trained a BERT-based large language model to answer questions about privacy policies in a Q&A format (not annotation) using a corpus of 1750 questions and 3500 expert answers. Other work that adapts LLMs for legal question-answering include (wan et al., 2023) and (Yue et al., 2023).
However, others have used LLMs specifically for privacy policy analysis. Tang et al., used various LLMs to annotate privacy policies via prompting  (Tang et al., 2023). However, the annotations are very simple (e.g., ”1st Party Collection”) and could likely be found using regexes. In contrast, our annotation approach uses a nuanced and theoretically grounded framework (GKC-CI) for normative privacy analysis. We believe LLMs are particularly suited for nuanced annotations as opposed to classical neural-network methods. This is because of LLMs’ increased representational capacity as well as their training on internet webscrapes, which exposes LLMs to cultural norms (Vaswani et al., 2017; Radford and Narasimhan, 2018; Devlin et al., 2018). However, we cannot provide a comparison to other researcher’s classical approaches for automated GKC-CI annotation because there are none developed in this space. To address this, we train an RNN on this GKC-CI task in our work.
3. GKC-CI Theory
The theory of contextual integrity (CI) (Nissenbaum, 2009) defines privacy as the adherence of information transfers, or “flows,” to sociocultural norms in specific contexts. For example, an information flow that might be appropriate between a patient and a doctor in a medical context (e.g., about a sensitive diagnosis) might not be appropriate between that doctor and their acquaintance in a recreational context.
CI further defines information flows as consisting of five essential parameters: 1) the sender of the information, 2) the recipient of the information, 3) the subject of the information, 4) the information content or attribute, and 5) the transmission principle that describes how or why the information flow occurs. The CI parameter annotation task entails identifying and labeling these five parameters in descriptions of information flows. For example, the CI annotation of: “We also collect contact information that you provide if you upload, sync or import this information from a device,” would label “we” as a recipient, “contact information” as an attribute, “you” as the sender, and “if you upload, sync or import this information from a device” as a transmission principle (example from (Shvartzshnaider et al., 2019a)).
Privacy Policy Sentence Parameter Annotated Text Parameter Annotated Text
If you consent, we may share information about Aim to provide…our partners Modality may
you with companies that aggregate it Attribute information Recipient companies that aggregate it
to provide analytics and measurement Condition If you consent Sender we
reports to our partners. Consequence N/A Subject you
Table 1. Example GKC-CI annotation. Italics and underlining added for emphasis.
The combined GKC-CI framework (Shvartzshnaider et al., 2022) further extends the CI framework, enabling the evaluation of strategies, norms, and rules drawn from theories of information governance. This allows cross-disciplinary research between the GKC and CI communities and allows a broader set of research questions to be addressed than either framework alone (Shvartzshnaider et al., 2022). Relevant to annotation, the combined framework divides the transmission principle into four categories drawn from the GKC institutional grammar: 1) aims and/or goals for specific actions, 2) conditions indicating when, where, or how aims apply, 3) modality operators implying pressure (deontics) or hedging, and 4) consequences, including sanctions for noncompliance, penalties in absence of consent, and benefits for proceeding. The GKC-CI parameter annotation task is identical to the CI annotation task except that it requires identifying the eight GKC-CI parameters instead of the five CI parameters. GKC-CI annotations thus provide more nuance than CI annotations at the expense of increased annotation difficulty. An example of what different GKC-CI parameters are present in a sample sentence is shown in Table 1. We encourage readers interested in additional information about the GKC-CI framework to refer to the original paper (Shvartzshnaider et al., 2022), which explains the theoretical foundations in detail and provides a worked demonstration of GKC-CI analysis in the Internet of things context.
4. Methods
We describe how we train LLMs to perform GKC-CI parameter annotation for privacy policies as accurately as possible. In doing so, we contribute the following insights.
First, we remark on how different LLMs of various architectures and sizes effect performance through ablation experiments. Because there is no other related work demonstrating how NLP models may be applied for GKC-CI annotation, we believe these insights may be particularly helpful to those considering normative privacy policy analysis in this space. Details of the model training process are provided in Section 4.4 and performance details are provided in Section 5. After evaluating a broad selection of models to find the most suitable candidate, we further tune the hyperparameters of our best performing model.
Second, we qualitatively examine the errors made by our best performing model to gain deep insights over the model’s behavior, particularly as it relates to practical deployment. We similarly believe that our approach to error analysis may provide a useful foundation for how others may analyze a model’s behavior in this space. The results of our error analysis are reported in Section 5.4.
Finally, we again use our best performing model to annotate a large set of privacy policies (Amos et al., 2021), which we report in Section 7, providing examples of the types of longitudinal and cross-industry analyses that can be performed via automated GKC-CI annotation.
4.1. Training and Testing Data
Our ground-truth labels are obtained by manually annotating GKC-CI parameters in 16 privacy policies from popular online services and e-learning websites, the exact breakdown of which is shown in Table 3 of Appendix A. We downloaded these privacy policies in HTML format and converted them to plain text for annotation. We used a customized version of the Brat Rapid Annotation Tool (Stenetorp et al., 2012) to manually label all GKC-CI parameters in the policies. In order to achieve consistent annotations across all annotators, we used a fixed set of guidelines defining each of the GKC-CI parameters (Appendix B). These guidelines were taken from (Shvartzshnaider et al., 2019a) for CI parameters and (Shvartzshnaider et al., 2022) for GKC-CI parameters to ensure continuity with prior work. Our ground-truth annotations included 6781 GKC-CI parameters across all 16 policies (Table 3). This ground-truth annotation process took two research assistants one semester to perform, including time spent learning the task.
In the process of annotating, we encountered several of the challenges discussed in (Shvartzshnaider et al., 2019a), including implicit parameters, ambiguous parameters, and policies not written with the CI framework in mind. We addressed these issues consistently with (Shvartzshnaider et al., 2019a). In general, the annotators made best judgment calls when faced with ambiguous parameters or difficult logic, consulting with the authors to ensure consistency. Importantly, we did not expect these manual annotations to be perfect. Rather, we treated them as best-effort annotations by researchers familiar with the task.
4.2. Formatting Examples
In formatting our examples, we wanted to ensure that the model is able to find the relevant text to be annotated given the presence of a parameter. Note that this necessitates teaching the model to filter through irrelevant text as well. As such, we include positive examples (examples where a parameter is present) and negative examples (text where a given parameter is not present, i.e., irrelevant text). The inclusion of negative examples is necessary to ensure that the model is usable in a real-world environment. Specifically, not every sentence of a privacy policy will include a GKC-CI parameter. By including negative examples, we ensure that our model will only output a parameter if one is present in its input sentence. We develop negative examples by taking positive examples and changing the parameter tag to one which does not occur in the sentence. We create a negative example for all possible negative tags. An example of a negative example is shown in the text immediately preceding Section  4.2.1.
We lightly format each sentence of a privacy policy as the basis of an example. We use sentences as the atomic unit for model input because sentence divisions are natural delimiters and previous work has shown reasonable accuracy with sentence-based annotation  (Wilson et al., 2018). We observe that other natural units of separation, such as paragraphs, tended to be particularly long due to the legal nature of privacy policies. This length ultimately resulted in problems where the paragraphs were too long to fit into the context window of some of our LLMs. [^4] Further, to our knowledge resolving a finite-length context window with large quantities of text is an open problem in the NLP space, although there has been substantial work in the area such as  (Press et al., 2022) and  (Beltagy et al., 2020). To ensure that training data was held consistent between LLMs, and because of the prior work above, we use sentences as the basis of our annotations. However, we also perform supplemental analysis with GPT-3.5 to verify that we are not critically disadvantaging our model by only feeding in sentences. The results of this are shown in Appendix  C.
Specifically, each example consists of the following parts: (1) a prefix to orient the LLM to the task, (2) a sentence from a privacy policy, (3) the GKC-CI parameter of interest, and (4) text delimiters. We include the text delimiters because modern LLMs decide what text to generate next based on all the text in their context window. As such, they cannot by default determine what text has been provided via the prompt, and what the LLM has generated. We choose the extremely minimal prefix “Annotate: ” for our task. We chose this minimal prefix to minimize the effects of prompt choice while still leveraging the training benefits of using a prompt  (Webson and Pavlick, 2022; Scao and Rush, 2021). We chose our text delimiters based on recommendations in OpenAI’s documentation, namely “–¿” and “x-x-x” respectively. Examples of how we fully format our examples are shown below:
  • Annotate: [“We also collect contact information that you provide”] Recipient–¿ Recipient: [“We”]x-x-x
  • Annotate: [“We also collect contact information that you provide”] Aim–¿ Aim: N/Ax-x-x
4.2.1. OpenAI models
Some chatbot models under OpenAI take an additional prompt, a system message, as input. GPT-3.5 Turbo is one such model. For models which take a system message, we replace the default system message from “You are a helpful assistant.” to “You are an assistant that understands Helen Nissenbaum’s theory of Contextual integrity (CI) and the governance of knowledge commons framework (GKC). This framework is abbreviated as GKC-CI. You reply with brief, to-the-point answers with no elaboration.” We do so to help orient the model to our task.
Finally, we perform an experiment leveraging the fact that Turbo is designed to respond to conversational inputs, as it is a chatbot. Specifically, we change our prefix from Annotate: to “For the following excerpt, provide the GKC-CI annotation of ’¡parameter¿’: ”. We call the model produced under this intervention GPT-3.5 Turbo, Prompt Engineered. Note that this is distinct from our baseline models that we prompt without fine-tuning.
4.3. Model Selection
4.3.1. Baseline Models
We consider two baselines for comparison against our fine-tuned LLMs: 1) a recurrent neural network (RNN) representing classical NLP approaches and 2) prompted non-fine-tuned LLMs.
For the RNN, we define BOS and EOS tokens, which is standard practice for these models (Robertson, 2024). We train the RNN in the exact same way as our fine-tuned LLMs, described in Sections 4.4.
For the prompted non-fine-tuned LLMs, we use GPT-4, GPT-4 Turbo, and GPT 3.5 Turbo. We use a fixed prompt format where we clearly delineate instructions from other text by placing instructions within brackets, thereby providing a structured framework that helps the model distinguish between the task directives and supplementary information. We also employ n-shot learning, a common prompting technique which has been shown to improve performance as n, the number of examples, increases  (Zamfirescu-Pereira et al., 2023; Brown et al., 2020; Gao et al., 2021). We use n={1, 3, 5} where the examples are chosen randomly from the training set.
4.3.2. Fine-Tuned Models
Model Family Open-Source Size(s) Considered Architecture Instruction-Finetuned RLHF Chat Variant
GPT-2 Yes Base (124M), XL (1.5B) Decoder-Only No No No
Flan-T5 Yes Base (248M), Large (783M) Encoder-Decoder Yes No No
Llama2 Yes 7B Decoder-Only No No No
GPT-3 No ”Davinci” Decoder-Only Unconfirmed Unconfirmed No
GPT-3.5 Turbo No Unknown Decoder-Only Unconfirmed Unconfirmed Yes
Table 2. A summary of the models we trained and their model families.
We consider five model families of diverse size and architecture: Flan-T5, GPT-2, Llama 2, GPT-3, GPT-4, and GPT-3.5 Turbo (Chung et al., 2022; Radford et al., 2019; Touvron et al., 2023b; Brown et al., 2020). Their properties are summarized in Table  2. In selecting models from these families, we wanted to choose from a wide range of high-performing or particularly usable LLMs. Approximately half of the models are open-source, while the GPT-3 models are proprietary. While the sizes of the GPT-3 models are not publicly released, GPT-3 and GPT-3.5 are likely the largest models we consider (their exact sizes are not public). We omit GPT-4 from our analysis because at the time of writing, the fine-tuning API is experimental. We also omit Llama2’s chat version from our analysis because, at the time of writing, it does not appear that Meta intended it to be fine-tuned further based on its GitHub repository (Facebook Research, 2023).
We now give a quick summary of how the various architectural features of the models we trained may potentially impact performance on the annotation task. First, we vary model size. Specifically, we consider models both in their “base” or default size [^5] When loading from the HuggingFace model hub as well as a larger size, if permissible by our hardware [^6] We employ an NVIDIA A100 for training, which we believe to be a reasonable baseline for the amount of compute other researchers will have access to.. We consider model size because the number of parameters in a model plays a large role in its performance (Hoffmann et al., 2022; Kaplan et al., 2020; Touvron et al., 2023a; Chung et al., 2022). However, smaller models may approach larger models’ performance if they are given more (pre)training data. We consequently include the Llama2 model to serve as an open-source proxy for larger models’ (e.g. GPT-3) performance  (Touvron et al., 2023a, b).
Most newer LLMs employ a “decoder-only” architecture, a deviance of the original design of the Transformer (Vaswani et al., 2017). Such models are referred to as encoder-decoder models, while models lacking an encoder block are decoder-only. Decoder only models can only ”look” at the preceding left tokens to determine what should be generated next, and tend to excel at creative or free-form generation  (Holtzman et al., 2020). In contrast, Encoder-Decoder models look at the entire sequence to determine what should be generated. Encoder-decoder models, like Flan-T5 or BERT, tend to perform well on tasks where the output is highly scoped by the input (Holtzman et al., 2020; Chung et al., 2022; Devlin et al., 2018). We include both Decoder-only and Encoder-Decoder models in our analysis because it is unclear what advantage, if any, an encoder-decoder model may have when annotating policies.
There exist a number of training paradigms which result in a model becoming more aligned to human intent. By alignment, we specifically say that a model is aligned if it produces outputs which are consistent with its human operator’s desires (assuming the model is capable of producing those outputs.)  (Chen et al., 2021; Ouyang et al., 2022). Because alignment broadly reflects a models’ ability to output text consistent with input tasks, we hypothesize that models which have undergone specific alignment training may perform better. We specifically look at two alignment mechanisms: Instruction-Finetuning and Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Wei et al., 2022). We include models in our analysis which are either confirmed or likely to have been trained according to these techniques.
Finally, we note that some models have been released as chat-models. This is relevant because 1) such models can be prompted in a different format from other non-chat models, and 2) these models are generally newer. While we are loath to conflate model age with performance, it is at least in the case of GPT and Llama that newer releases tend to eclipse older releases (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b). Thus, along with reasons explained in section 4.4, we include the fine-tunable chat model GPT-3.5 Turbo in our analysis.
4.4. Model Training
All 16 manually annotated privacy policies used for model training and testing underwent processing as outlined in section 4.2. We randomly reserved 70% of the manual annotations to constitute our training data (21,588 examples), while the other 30% (9252 examples) are reserved as testing examples. We specifically use low-rank adaption (LoRA), a type of parameter efficient fine-tuning (PEFT), as our training method instead of traditional fine-tuning to ensure our open-source models are being trained in a way which is consistent with proprietary models (Hu et al., 2021). PEFT methods in general encompass methods of training some of a model’s parameters, but not all (i.e., being ”parameter efficient” when training). LoRA describes a method of being parameter efficient when training by freezing a model’s weights and inserting trainable rank-decomposition matrices into the model’s architecture at each layer (typically, the attention layer). Then, the low-rank matrices are ”trained” during fine-tuning, while the base model remains frozen. We use LoRA because OpenAI’s business model suggests that LoRA is being employed in the place of traditional fine-tuning, as low rank matrices are very cheap to store and entire models are typically large and expensive. [^7] We make this claim because OpenAI allows for “fine-tuning” through their API. Traditional fine-tuning would require making full copies of the model (e.g., GPT-3.5) for each user. Doing so would result in terabytes of space being allocated per user due to the size of OpenAI’s models. Additionally, it has been observed that previously fine-tuned models under OpenAI may change in performance without warning, as OpenAI routinely updates their models. This behavior would not be observed if traditional fine-tuning were occurring because each user would have their own discrete copy of the model.
Thus, to ensure that all our model comparisons are apples-to-apples comparisons, we ensure that all models are trained using LoRA. We keep the training parameters constant between all open-source models. We make an exception for the learning rate of Flan-T5, as the model’s documentation recommends a slightly higher learning rate than the other models. This higher learning rate could cause the model to converge faster than others.
In training, we perform a number of ablation experiments, changing the following in a full-factorial manner for the open-source models:
  1. (1)
    Training Epochs, where we vary the number of epochs the model trains for. Training a model for a single epoch means that the LLM sees every example in the training set once. Training for n epochs means that the LLM sees every example in the training set n times. Varying the number of training epochs thus influences the quantity of training tokens the model is exposed to, as well as the number of repeated examples the model is exposed to.
  2. (2)
    Example formatting, where some examples are prepended and appended using the model’s BOS, EOS tokens. BOS (Beginning of Sentence) and EOS (End of Sentence) are typically defined and used during LLM pre-training to semantically capture the start and end of a training document. We hypothesize that the inclusion or exclusion of these tokens during fine-tuning may have a subtle effect on the model’s performance.
We note that Open-AI’s proprietary models do not offer the same number of training options as the open-source models, and are ultimately unable to perform the example formatting experiment with Open-AI’s models. We ultimately benchmark 50 models distributed between Open-AI’s and our open-source models.
5. Model Performance
5.1. Metrics
We begin a discussion of performance by clearly defining our performance metrics. Specifically, our main metric is accuracy. That is, we consider the model’s output to be correct or accurate if it is an exact string match of the ground-truth. We do not directly consider precision and recall because they are less intuitive for this task. We envision that downstream users are likely to be particularly interested in the performance per GKC-CI parameter. As such, we focus on per-parameter accuracy.
As we want to capture nuanced model errors, we categorize each model output into exactly one of four possible results: perfect match, superset match, match error, or identification error. These categories are defined as follows:
  1. (1)
    Perfect Match indicates the model’s annotation is an exact string match with that of the human annotator. Only the Perfect Match represents a correct annotation.
  2. (2)
    Superset Match indicates the model’s annotation contains all the words of the human annotator. However, the model may have highlighted additional information or may have included information which does not appear as a contiguous lump of text in the policy.
  3. (3)
    Match Error captures instances in which the model agreed that a certain parameter was present, but did not identify the “correct” annotation. This can include completions that are flat-out incorrect, completions that don’t identify the correct number of instances of a parameter in the input, and completions that have identified a proper subset ($\subset$) of the correct words.
  4. (4)
    Identification Error occurs when the model, despite being prompted with a specific parameter (e.g., “Aim”), failed to include that parameter in its completion.
Because any model output is categorized as belonging to one of the above categories, accuracy is defined as:
Accuracy =Correct OutputsAll $\text{Accuracy }=\frac{\text{Correct Outputs}}{\text{All Outputs}}=\frac{PM}{PM+SM+ME+IE}$
where $PM$ is the number of perfect match annotations, $SM$ is the number of superset matches, $ME$ is the number of match errors, and $IE$ is the number of identification errors.
5.2. Baseline Results
Our baseline RNN model had a test set accuracy of only 6% (Figure 1). While the RNN appropriately handled negative examples, it struggled with positive examples. Specifically, the RNN failed to accurately annotate any positive examples, while all LLMs other than GPT-2 were able to successfully annotate at least one positive example.
The prompted non-fine-tuned LLMs performed even worse, with the best such model having a test set accuracy of only 2% (Appendix D). While more effective prompting strategies are likely to exist, we believe that this finding provides evidence that LLMs should not be applied off-the-shelf for GKC-CI annotation as of now.
Comparing these results to the performance of the fine-tuned LLMs below, it is evident that fine-tuned LLMs substantially outperform both classical NLP RNNs and non-fined-tuned LLMs. This justifies the time and expense needed for fine-tuning and further indicates the non-triviality of the GKC-CI annotation task.
5.3. Fine-Tuned Results
We benchmark each of our 50 models on each of the 9252 sentences in our test set. The results of our benchmark are shown in Figure 1. We find that OpenAI’s proprietary models perform significantly better than any of the other open-source models we test. Additionally, no open-source model that we considered had performance high enough to be considered even a poor substitute. Examination of the open-source model’s outputs indicates a lack of substantial training convergence within 10 epochs for all models. Graphs of the open source models’ performance at 1, 5, and 10 epochs are included in Appendix E. We comment that each of the open-source models took ¡12 hours to train, but that training time is extremely sensitive to specific hardware configurations. We hypothesize that given enough compute, the open-source models may eventually reach the performance of Open AI’s proprietary models. We briefly summarize some notable lessons from working with the open-source models, then turn our discussion to the OpenAI models.
Figure 1. Test set performance of the top-performing models variants, including the RNN, with $\leq$ 10 epoch of training. GPT3,5_TPE refers to the prompt-engineered version of GPT-3.5 Turbo, GPT3,5_TG refers the generic GPT-3.5 Turbo model, and GPT3,5_t2s refers to the joint performance of the GPT-3.5 Turbo, 2-Step models.
5.3.1. Open-Source Takeaways
Firstly, we note that model size does play a significant role in model ability in relation to our annotation task. GPT-2 and Flan-T5 both failed utterly at their smaller model sizes, while their XL and Large variants performed $\sim$ 3% to 15% better. We additionally note that absolute model size in terms of parameters does not appear to be a consistent indicator of performance across model families, at least at the small sizes we consider. Llama2’s 7B variety appears to perform similarly to GPT2’s XL variety, despite being over four times the size. We mention this to caution resource-constrained researchers: larger may not always mean better when performing cross-model family comparisons. Additionally, those models which were aligned (Flan-T5 and all GPT3.5s) appeared to perform the best. We recommend researchers prioritize smaller, aligned models.
Secondly, we observe that models appear to be vulnerable to the inclusion or exclusion of BOS, EOS tokens. Namely, we observe small performance differences between all open-source models, but most strikingly in the Llama2 model family. We believe this is significant because it suggests that opaque defaults have observable effects on performance. Further, those effects do not appear to be consistent across model families. Specifically, as of the time of writing, HuggingFace’s tokenizers all have BOS, EOS tokens internally defined, but each model has a different default behavior when it comes to including or excluding BOS, EOS tokens. We urge other researchers to be particularly careful of library and model defaults as they could be a potential confound in model performance.
Figure 2. Per GKC-CI parameter performance for our best performing model, GPT 3.5TPE_25ep.
5.3.2. OpenAI Models
Appendix F summarizes model ability on positive examples, which are those examples that contain a GKC-CI parameter. We observe that models vary substantially in their per-parameter performance.
All three models are variants of GPT-3.5 Turbo. GPT3.5-Turbo, two step 10 epochs refers to a two model system where the first model determines if the text contains a GKC-CI parameter and the second model identifies the parameter. For this model, an example is a perfect match if and only if the first model and second model classifies the prompt correctly. This model performed the worst out of the three at 30% correct positive examples and 97% correct negative examples. While the model demonstrates commendable success in identifying negative examples, its limited capacity to accurately identify positive examples significantly undermines the overall efficacy of its performance.
GPT3.5-Turbo, Generic 10 epochs and GPT3.5-Turbo, Prompt Engineered 10 epochs performed similarly well with GPT3.5-Turbo, Generic 10 epochs at 73% correct positive examples and 88% correct negative examples and GPT3.5-Turbo, Prompt Engineered 10 epochs (hereafter, GPT 3.5TPE_10ep) at a slightly better 75% correct positive examples and 89% correct negative examples.
Due to this strong performance, we further examine GPT 3.5TPE. We speculate that the reason why GPT 3.5TPE_10ep performs better than GPT 3.5TP_10ep is due to the added context from prompt engineering. Specifically, prompt engineering may provide enough context for GPT 3.5TPE_10ep such that the model is able to relate the annotation task to information it has been trained on. This phenomenon has been more broadly observed in the NLP space (Scao and Rush, 2021; Webson and Pavlick, 2022).
Like the open-source models, we similarly investigate how the number of training epochs effects the performance of this model. The results of this are shown in Figure  4. We observe that GPT 3.5TPE starts to have diminishing returns after 25 epochs worth of training. We hypothesize that at 50 epochs, it is likely that the model has overfit. We further investigate the effect that training data quantity has in Figure 4. We note that the model starts to have diminishing returns in terms of performance per data quantity at the 75% mark, and that 50% of our training data appears to be the necessary minimum for our model to begin to converge. Because we observe no detriment to using the full quantity of our training data, we do not restrict ourselves to a subset of our training data.
Finally, to ensure that our model’s performance isn’t critically hindered by being trained exclusively on sentences, we tested a fine-tuned version of GPT 3.5TPE, our most successful model, using paragraphs instead of sentences. The goal was to determine whether providing more context in the input would improve performance. Despite subjecting this model to the same epoch ablation experiments as our top-performing model, it only achieved a maximum accuracy of 67.75%. For a detailed comparison of the performance between both epoch ablation experiments, please refer to Appendix C.
To summarize, we observe that the model benefits from the full training data and that the model may be at risk of overfitting after 25 epochs. We consequently choose our top performing model to use the full training data, and train it for 25 epochs, which took 3 hours. We specifically call our top model GPT 3.5TPE_25ep, and its performance is shown in Figure  2. To better understand the strengths and weaknesses of this model, we perform qualitative error analysis on the model’s errors.
Figure 3. GPT 3.5TPE’s performance on the test set at 1, 5, 10, 25, and 50 epochs. Perfect Matches and Superset Matches were considered to be ”correct”.
Figure 4. GPT 3.5TPE_25ep’s accuracy on the test set as training data increases. Perfect Matches were ”correct.”
5.4. Qualitative Error Analysis
In order to better understand the errors made by GPT 3.5TPE_25ep, we performed qualitative coding on the 37 match errors for positive examples produced by the model, i.e., match errors where neither the ground truth label nor the model completion was “N/A”. This served two purposes.
First, we identified cases where labels were mistakenly labeled as errors, specifically where the model annotation was semantically equivalent, albeit syntactically different from the ground truth.
Second, we conducted a detailed analysis of the model’s performance, examining the factors contributing to both its strengths and weaknesses. This provided more confidence in overall model performance.
Identifying trends in these errors offers valuable insights into the model’s behavior. Although these insights may not explain why an error occurs (LLM models have notoriously poor explainability (Druce et al., 2021; Barredo Arrieta et al., 2020)), they help users become aware of the model’s limitations. This enables users to use the model more strategically, leveraging its strengths and compensating for its weaknesses, thereby enhancing its overall effectiveness.
5.4.1. Qualitative Coding
To ensure the reliability and consistency of the coding process, two expert coders initially met to collaboratively develop a comprehensive codebook consisting of ten codes: three parent codes and seven child codes. The full codebook can be seen in Appendix G.
After joint codebook creation, each coder independently coded all match errors produced by GPT 3.5TPE_25ep. Once the coding was complete, we computed inter-coder reliability and found a high level of agreement between the two coders with a Cohen’s kappa score of .87 (Cohen, 1960). The results of the qualitative coding are visualized in Figure 5 and detailed further below.
5.4.2. Semantic Equivalence
The code “Semantic Equivalence”, which is the sole code of the Semantic Equivalence category, was the most prevalent in our error analysis—accounting for 19/37 (51%) errors. The following two examples demonstrate what this error looks like. The text in quotation is the expert annotation, while the underlined sections are the model’s response:
  1. (1)
    Aim: “to help us operate or administer the Services
  2. (2)
    Recipient: “These Services
Note that the model’s response only differed by an article or an adjective, and both are equivalently correct annotations.
Figure 5. Breakdown by category of the various types of errors found from our qualitative analysis.
5.4.3. Incorrect Expert Labels
The category Expert Labels Is Wrong had nine examples (24%). Occasional expert mis-annotations are expected for a task of this complexity. We are encouraged that there were relatively few examples in this category, supporting the quality of our ground truth. Importantly, for the examples in the “Expansive Ground Truth” (4/9, 44%) and “Partial Ground Truth” (3/9, 33%) child codes, the model performed the task more correctly than the expert annotator—either by omitting superfluous words included in the expert annotation or including necessary ones the expert annotator missed.
Consider the following example “expansive ground truth” annotation. The text in quotation is the expert annotation, while the crossed sections are what the model correctly excluded.
  1. (1)
    Consequence: “You can set your browser to not accept cookies, but this may limit your ability to use the Services.”
Conversely, the following example “partial ground truth” annotations show the expert annotation in quotations, while the underlined words are what the model correctly choose to additionally include in its response.
  1. (1)
    Recipient: “trusted companies” that work with, or on behalf of, Crowdmark to process information
  2. (2)
    Condition: “to comply with its general obligations under the GDPR,” in particular to process the personal data it collects in accordance with Articles 5 and 6, and to comply with Articles 13, 14, 24, 30 and 32, and to comply with any actionable rights of the data subject
Combined, “Expansive Ground Truth” and “Partial Ground Truth” codes represent only 9 of the 37 errors (24%); however, we emphasize this finding as particularly exciting because they demonstrate that the model can identify precise annotations for the requested parameter. In other words, the model’s ability has surpassed that of our highly-trained human annotators for these situations.
5.4.4. True Model Errors
The category “Completion Errors” accounted for 9 out of 37 coded examples (24%). Notably, a significant majority of these errors (6 out of 9, 67%) fall under the “meaningful subset” child code. A “meaningful subset“ annotation included a segment of the correct response, but missed words that altered the meaning of the annotation. In the following example, the text in quotation is the expert annotation while the underlined sections are the model’s response:
  1. (1)
    Aim: “solely for the purposes of providing the relevant services to Kaltura
  2. (2)
    Attribute: “personal data, any communications or material of any kind that you e-mail, post, or transmit through the Site, such as questions, comments, suggestions, and other data
Observe that the model’s completion is a sentence fragment, regardless of the parameter specified. Many of these fragments do encompass enough of the correct answer for someone well-versed in CI-GKC to grasp the intended annotation. However, we consider these incomplete responses to be incorrect even though the model’s answer is a meaningful subset of the correct response.
Furthermore, the code “Completion is wrong” contains the remaining elements Completion Is Wrong category, comprising the final 3 out of 9 codes (33%). All of these responses labeled text from within the sentence with no relation to the actual GKC-CI parameter. For example the model incorrectly labels the underlined part of the following sentence as an attribute:
  1. (1)
    Attribute: “from the institution including the user’s identifier and organizational affiliation”
Our qualitative analysis offers a comprehensive view of the model’s performance, detailing both its strengths and limitations. Appreciably more than half of the purported match errors can be attributed to “Semantic Equivalence”, “Expansive Ground Truth”, and “Partial Ground Truth.” These combined child codes constitute 27 examples, resulting in a $0.29$ percent increase in the number of correct annotations overall. This implies that our benchmarking metrics for model accuracy serves as a conservative estimate of model performance. In essence, while our model displays commendable performance, there is still room for improvement to address model errors.
5.5. Training Data Review
Motivated by the nine expert labeling errors we uncovered in Section 5.4.3, we had two different CI experts review a random sample of 5% of our training data. These experts agreed with the training data on 90.57% of the sample, with the most common disagreement being the identification of occasional attribute parameters that were missed in the original annotations. The original annotators spent several months annotating thousands of parameters, so it is unsurprising that a few were missed. Furthermore, governing knowledge commons and contextual integrity are the topic of active research and subject to different interpretive nuances by different members of the community, which is reflected in different annotations. This training data review validates our results, as our best performing model’s 90.65% accuracy is close to the inter-expert annotation agreement, suggesting that the model approaches human performance on the task.
6. Limitations
Despite the high accuracy of our best performing model, we point out a few limitations. First, website privacy policies are not written with the GKC-CI framework in mind, so a perfect mapping from text to GKC-CI parameters may not be possible. We do not know the maximum performance that can be achieved by GKC-CI annotation, and it likely varies depending on the specific documents annotated. While we have created and used the largest training set for this task to date, and our analysis indicates that further increasing training set size would not substantially improve performance (Figure 4), it is possible that training data from a different set of policy documents might affect overall performance.
Second, the LLMs we use do not have information about the input text outside of the context window. This is unlike a human lawyer, who could review the entire document while identifying information flow descriptions. Fortunately, the information needed to identify GKC-CI parameters typically appears in the text immediately surrounding the parameter itself. The decreased accuracy we observed when using a longer context window (Section 5.3.2, Appendix C) supports our use of one sentence contexts.
Third, we do not perform coreference resolution, i.e., identifying when different sections of text refer to the same real world entity. For example, a sentence may refer to the website user multiple times, with each instance receiving a GKC-CI parameter annotation (e.g., “sender”). Although our model would accurately annotate each mention of the sender, it would not automatically recognize that these annotations refer to the same legal entity. If that were necessary for a downstream application, another model would be required. Since coreference resolution is an active research problem in its own right, and is independent from GKC-CI parameter labeling, we leave this for future development.
7. Example Applications
We applied our GPT 3.5TPE_25ep model to 456 privacy policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021) to demonstrate the type of analyses enabled by GKC-CI annotation at scale. This dataset contains over 1 million privacy policies from over 100,000 companies spanning more than two decades, making it an ideal data source. However, we note that the primary contribution of our project remains the LLM training and evaluation (Sections 45). This section is not meant to provide a comprehensive analysis of policies in the Princeton-Leuven dataset. Rather, we intend the following examples to inspire future work using our LLM annotation method and the annotated policies we provide. All 456 policies we annotated for the following analyses are publicly available at the GitHub repository for this paper. [^8] https://github.com/JakeC007/Automated_GKC-CI_Privacy_Policy_Annotations
7.1. Longitudinal Privacy Policy Analysis
First, we chose 8 prominent companies and organizations [^9] Facebook, The New York Times, GitHub, Buzzfeed, Google, Bank of America, Electronic Frontier Foundation (EFF), and the National Science Foundation (NSF) representing a variety of sectors, including “big tech,” news, entertainment, finance, and government. This variety is useful because each sector has a unique approach to data collection, user engagement, and compliance with privacy regulations. Furthermore, these websites have undergone varying levels of public scrutiny. For instance, while Facebook and Google have faced major privacy debates, leading to numerous changes in their privacy policies, entities like nsf.gov operate under distinct governmental standards. This list also highlights geographic diversity concerning headquarters and user base, with some companies primarily serving U.S. audiences, like Bank of America, while others have a global reach, necessitating compliance with various international privacy laws per the Brussels effect (Bradford, 2012).
For each of these companies and organizations, we used our model to annotate one privacy policy from every year that the company or organization appears in the dataset (128 policies total). The number of parameters in the policies of each company or organization over time are presented in Appendix H (Figure 17 and Tables 56).
The results provide insight into the evolution of privacy policies. For instance, we notice a generally increasing trend in the number of GKC-CI parameters included in privacy policies over time. As specific examples, the privacy policies of Buzzfeed and GitHub described fewer than 60 GKC-CI parameters in their policies from 2008-2010, but now describe over 400 or 500 parameters, respectively. The increase in parameters in the GitHub policy from 2016 to 2019 corresponds to the acquisition of GitHub by Microsoft.
The EFF and the NSF show similar, if less dramatic, increases in the number of parameters over time. This trend mirrors previously documented increases in average privacy policy length from 1996 to 2021 (Wagner, 2022), providing a sanity check for our method – we expect longer privacy policies to include more details about information transfers. Indeed, the New York Times privacy policy underwent a dramatic decrease in the number of GKC-CI parameters in 2006, corresponding to an approximately 80% decrease in the length of the policy (23736 to 4912 words). The number of parameters then increased to above its previous maximum in 2011 when the policy length increased to 33616 words.
We also notice that although parameter counts are generally increasing and roughly track total policy length, the relative percentages of different parameters remains nearly consistent within each company’s policy. This suggests that although companies are adding additional details to descriptions of data transfers in their privacy policies, these additions are not broadly skewed toward specific parameters. To understand the importance of this result, consider some counterfactual examples: If the relative percentage of aim parameters were to have increased, it would indicate that organizations are increasingly using privacy policies to inform why information is being collected over what information is being collected. If the relative percentage of attribute parameters were to have increased, it might indicate that organizations are collecting more data types per information transfer. While we don’t see either of these trends for these 8 companies, we examine the variance of parameter types in the privacy policies of a larger set of companies in Section 7.2.
A closer look at the GitHub privacy policies from 2016 and 2019 demonstrates how GKC-CI annotation can be used to automatically identify privacy policy updates of interest. Specifically, the 2019 policy includes more aim parameters describing new ways that the site may use collected information, such as “to make recommendations for you, such as to suggest projects you may want to follow or contribute to” and “to determine your coding interests.” Noticing a substantial increase in the number of aims and focusing directly on those new parameters allows for quick assessment of changes in information use. Similarly, automatically identifying new or changed attributes would allow quick assessment of new types of information collected.
Privacy policy updates happen regularly on a vast number of online services, and it is difficult for experts, to say nothing of average users, to identify which updates are salient for particular privacy concerns. We anticipate that our LLM annotation approach could be used in the future to automatically sort updates by parameter type and, combined with a visual interface, faster review of privacy policy updates.
7.2. Cross-Industry Privacy Policy Analysis
We next used our fine-tuned LLM to annotate the most recent privacy policies within the Tranco top 300 (Pochat et al., 2018) websites that are in the Princeton-Leuven corpus (164 policies total).
In each of the following analyses, we highlight extreme examples from across these 164 policies to demonstrate how annotation at scale facilitates directed data exploration. Previous work has shown that detailed analysis of individual annotated policies using the CI framework can identify specific ambiguities and normative shortcomings (Shvartzshnaider et al., 2019a). While deep analysis of individual policies is out of scope for this paper, the following paragraphs show how GKC-CI annotation can be used to identify policies that might be worth such detailed exploration in future work.
Parameter type variance.
We first calculated the variance in the percentages of individual parameter types across all annotated parameters in each policy. Previous work using CI annotation emphasized that descriptions of information transfers that are missing specific parameter types or that included substantially more specific parameter types (“parameter bloating”) lead to ambiguities about the actual data handling practices of the organization (Shvartzshnaider et al., 2019a). Since policies with a greater variance in the percentages of individual parameter types are more likely to exhibit these issues, we rank our annotated policies by this metric.
Figure 6 shows the fifteen policies with the highest variance of parameter type percentages. This includes a policy from apache.org with relatively few aim parameters and a policy from mozilla.org with relatively few attribute parameters. A quick review of the mozilla.org policy shows why – it defines all data covered by the policy in one sentence: “For us, ‘personal information’ means information which identifies you, like your name or email address.” The policy does not say which particular forms of personal information are connected to the descriptions of information use in the rest of the policy, leaving it ambiguous whether e.g., what specific customer data might be shared with other entities for “processing or providing products and services.” While this type of high-level analysis doesn’t necessarily imply the existence of policy ambiguities, it suggests that policies with high parameter type variance are promising candidates for a detailed evaluation through the lens of the GKC-CI framework. The parameter percentages for all 164 privacy policies are provided in Tables 78 in Appendix I.
Figure 6. The 15 privacy policies with the highest variance in the percentage of individual parameter types across all parameters annotated in the policy.
Parameter to sentence ratio.
We next calculated the ratio of annotated GKC-CI parameters to the number of sentences in each policy. This provides a metric of the “density” of information transfer descriptions in the policy. Figure 7 shows these data for the 15 privacy policies with the highest ratio of annotated parameters to sentences. The ratios for all 164 privacy policies are provided in Tables 910 in Appendix J. The top 15 policies include those from content distribution networks (b-cdn.net), web component frameworks (ampproject.org), Facebook, Microsoft (bing.com), Apple (icloud.com), Zoom, and social media websites (linkedin.com, snapchat.com, t.co, tumblr.com), among others.
While these policies may exhibit parameter bloating issues due to the density of parameters, they may also be good examples of policies providing meaningful details about data handling practices. The first few policies on this list provide a microcosm of this variety. The b-cdn.net policy is very minimal, but each sentence is a short, to-the-point data handling description. The Facebook policy is longer, but an earlier 2018 version was identified in (Shvartzshnaider et al., 2019a) as having parameter bloat. Either way, directing future in-depth investigations toward policies with high parameter density would provide examples of GKC-CI information flow descriptions for case studies for teaching (Apthorpe, 2021) or iteration on the GKC-CI framework.
Figure 7. The 15 privacy policies with the highest ratio of GKC-CI parameters to sentences out of all 164 of the Tranco top 300 websites in the Princeton-Leuven corpus.
Website popularity.
As a comparison against the 164 popular websites from the previous analyses, we next annotated the 164 policies from the Princeton-Leuven corpus with the lowest Tranco rankings (999981 to 991993). Figure 8 compares the distributions of total parameter counts and parameter to sentence ratios between these sets of websites.
The distributions of total parameter counts are significantly different ($p<0.005$, Mann-Whitney U test), with the more popular websites having more GKC-CI parameters in their privacy policies on average than the less popular websites (mean $412$ versus $130$). This result makes sense, as more popular websites are under more scrutiny about their handling of user information and therefore include more information about data practices in their privacy policies. However, the distributions of parameter to sentence ratios are not significantly different ($p>0.05$, Mann-Whitney U test), indicating that the more popular websites are generally providing more details by adding to policy length, rather than by increasing the density of information flow descriptions.
We anticipate that automated GKC-CI parameter annotation will enable more detailed statistical analyses across privacy policies in the future. For example, one could track the changing correlation between condition parameters in privacy policies and the text of data privacy regulations as both policies and regulation are updated – identifying outliers for scrutiny.
Figure 8. Comparison of all 164 websites in the Tranco top 300 versus the 164 websites lowest on the Tranco rankings in the Princeton-Leuven corpus. Top: Distribution of the raw number of GKC-CI parameters annotated by our model. Outliers with $>700$ parameters (13 policies for the more popular websites and 1 policy for the less popular websites) are omitted. Bottom: Distribution of the ratio of annotated GKC-CI parameters to sentence count.
7.3. GKC-CI Annotation Visualizer
We have developed a visualizer tool designed to enhance the interpretability of our LLM’s privacy policy annotations. This tool was created to empower users to better understand the complex data flows and governing principles that underpin privacy policies by visually distinguishing between different parameter tags. By highlighting various aspects of the policy in different colors, the visualizer makes the described information flows more accessible, allowing users to quickly grasp how their data may be used and protected under the terms of the policy.
The tool functions as a local Python script that processes both the raw text of the privacy policy and the output log from the LLM. It begins by matching specific text segments from the policy with those used in the LLM prompts. The visualizer then highlights these segments in different colors based on the associated annotations, making it easy for users to see which parts of the text correspond to specific GKC-CI parameters. This highlighted text is presented in a user-friendly GUI window, where users can scroll through and interact with the augmented policy. Additionally, it generates an output file that explains the highlighted segments, helping users understand the implications of the annotations.
To illustrate the capabilities of the visualizer, we applied it to two versions of Facebook’s privacy policy from 2015 and 2019. Comparing visualized excerpts of these versions (Appendix K) effectively highlights how data practices and governing principles have evolved. For instance, there are more GKC-CI parameters in the 2019 excerpt than the 2015 excerpt, particularly more aims and consequences. The new policy does a better job informing users about the repercussions of the described data collection. The visualizer’s color-coded presentation makes it easier to understand how the policy has changed and will facilitate future research using GKC-CI annotation.
8. Future Directions
As demonstrated in Section 7, the ability to accurately annotate privacy policies with GKC-CI parameter tags enables a variety of previously infeasible analyses. We are excited about the potential for the GKC and CI community to use our method to facilitate advances at the intersection of machine learning and privacy.
A common question posed by CI researchers is whether information flows (composed of GKC-CI parameters) are appropriate in their respective contexts. This is a core element of the CI framework, which understands privacy breaches as inappropriate information flows that violate contextual norms (Nissenbaum, 2009). There are many ways to discover the norms of a particular context, including surveys (Apthorpe et al., 2018, 2019; Shvartzshnaider et al., 2016), interviews (Kumar et al., 2024), focus groups (Shaffer, 2021), and textual analysis (Barth et al., 2006), among others. We propose using LLM annotation as the first step in a multi-method pipeline for this task: 1) a researcher extracts GKC-CI parameters from segments of an organization’s privacy policy using LLM annotation, 2) the parameters are piped into a survey template about contextual appropriateness (like that in (Apthorpe et al., 2018)), 3) crowdsourced survey responses from relevant community members indicate (mis)alignment with existing norms. This entire process could be automated from the perspective of the researchers, enabling much larger scale CI-based audits of data handling practices across many organizations than previous possible.
We also believe that GKC-CI annotation with LLMs need not be limited to privacy policies. Many documents, from white papers to media reports, describe transfers of information. We expect that fine-tuned LLMs will be able to identify GKC-CI parameters as accurately in those documents as in privacy policies. This opens an new set of applications that we hope the community will pursue.
9. Conclusion
This paper demonstrates that high-accuracy annotation of contextual integrity (CI) and governing knowledge commons (GKC) parameters in privacy policies can be achieved using LLMs. We ultimately find that GPT 3.5TPE_25ep had the best performance, with an accuracy of 90.65% for exact string matches. While we find that the proprietary LLMs outperformed open-source models, we report some valuable findings for researchers interested in performing LLM application studies. Namely, 1) that LLM size must be considered in context to model family; smaller, aligned models be the most economical choice, and 2) that library defaults are likely to introduce confounds and should be checked.
We demonstrate the usefulness of our fine-tuned model by annotating 456 privacy policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). We demonstrate that large-scale GKC-CI annotation can be an effective tool for data exploration, highlighting changes in parameter frequency over time, policies with relatively high variances across parameter type percentages, and policies with relatively high ratios of parameters to sentences. This facilitates automated review of privacy policy updates to identify meaningful changes in information flow descriptions, potentially with normative implications. We have made our model training code, training and testing data, annotation visualizer, and all annotated policies publicly available [^10] https://github.com/JakeC007/Automated_GKC-CI_Privacy_Policy_Annotations in the hope that this work motivates future use of GKC-CI parameter annotation.
Acknowledgements
We thank research assistants Sophia Goffe and Wael Mohamed for their contributions. This research was funded by the Colgate University Faculty Research Council and the Colgate University Department of Computer Science. Some of the results presented in this paper were obtained using the Chameleon testbed (Keahey et al., 2020) supported by the National Science Foundation.
Appendix A Ground Truth Details
Company Word Count Sender Subject Recipient Attribute Aim Condition Modality Consequence
Cengage (Cengage, 2022) 3340 12 15 34 44 64 82 30 0
Crowdmark (Crowdmark, 2022) 3216 12 28 37 42 80 92 42 6
Dropbox (Dropbox, 2022) 2485 10 9 20 26 36 30 10 4
Facebook (Facebook, 2022) 4151 40 48 74 113 78 84 42 0
Gradescope (Gradescope, 2022) 11431 18 20 39 69 104 152 28 2
Honorlock (Honorlock, 2022) 1199 10 9 11 22 18 32 22 2
Kultura (Kultura2022, 2022) 6255 15 29 54 63 96 164 52 4
LinkedIn (LinkedIn, 2022) 6298 37 58 80 111 110 174 22 0
Matlab (Matlab, 2022) 5580 27 27 61 85 98 150 44 4
Niantic (Niantic, 2022) 5539 27 33 44 63 92 94 16 2
NYTimes (The New York Times, 2022) 5000 12 25 41 50 50 82 18 2
Packback (Packback, 2022) $4$ 11 14 25 35 40 94 18 4
Panopto (Panopto, 2022) 4167 17 16 28 34 62 82 42 2
Proctorio (Proctorio, 2022) 9353 28 24 61 89 124 140 54 2
Stripe (Stripe, 2022) 7460 38 48 73 96 110 122 40 4
Turnitin (Turnitin, 2022) 10220 15 24 24 52 94 90 18 4
Table 3. Number of labeled parameters in ground-truth GKC-CI annotations of 16 privacy policies from popular websites and e-learning services.
Appendix B Brat Annotation Legend
Figure 9. Legend in the customized Brat Annotation tool for expert annotators to use as reference.
Appendix C Performance On Paragraph Delineated Data As Epochs Vary
Figure 10. The dotted orange line (“FT-Paragraph Accuracy”) is the test-set accuracy of GPT 3.5TPE trained on paragraph delineated data at 1, 5, 10, 25, and 50 epochs. The solid blue line (“FT-Sentence Accuracy”) is the test-set accuracy of GPT 3.5TPE trained on sentence delineated data at 1, 5, 10, 25, and 50 epochs. An accurate annotation is one that is an exact string match.
Appendix D Baseline Non-fine-tuned performance With N-Shot Learning as N varies
Figure 11. Test set accuracy of non-fine-tuned GPT 3.5TPE with simple prompting and 1, 3 and 5 few shot examples. An accurate annotation is one that is an exact string match.
Appendix E Open-Source Model Performance as Epochs Vary
Figure 12. Test-set accuracy of our RNN at 1, 5, and 10 epochs. An accurate annotation is one that is an exact string match.
Figure 13. Test-set accuracy of Flan-T5 models at 1, 5, and 10 epochs. An accurate annotation is one that is an exact string match.
Figure 14. Test-set accuracy of GPT-2 models at 1, 5, and 10 epochs. An accurate annotation is one that is an exact string match.
Figure 15. Test-set accuracy of Llama2 models at 1, 5, and 10 epochs. An accurate annotation is one that is an exact string match.
Appendix F Performance Breakdown For The Three Top Performing Models
Figure 16. A comparison of our top-performing models on positive examples by GKC-CI parameter. GPT3,5_TG_10ep refers the generic GPT-3.5 Turbo model, GPT3,5_TPE_10ep refers to the prompt-engineered version of GPT-3.5 Turbo, and GPT3,5_T2S_10ep refers to the two model GPT-3.5 Turbo system. All models were trained for 10 epochs.
Appendix G Codebook
Code Description
Completion Errors
Completion Is Wrong Completion is outright incorrect failing to provide the accurate answer.
Meaningful Subset Completion partially captures the correct response but falls short of completeness.
Completion Over-labeled Completion includes correct answers but erroneously incorporates nearby words into the parameter tag.
Expert Labeling Errors
Expert Labels Is Wrong Expert label itself is incorrect.
Expansive Ground Truth Expert label is correct but overly broad and the completion offers a more precise response.
Partial Ground Truth Expert label misses a portion of the correct label, but the completion captures it accurately.
Semantic Equivalence
Semantic Equivalence Completion and the ground truth label differ in wording but convey equivalent semantic meanings.
Table 4. Codebook for qualitative error analysis. Parent codes in bold.
Appendix H Longitudinal Data
Figure 17. Number of annotated GKC-CI parameters in the privacy policies of 8 prominent companies over time. Policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). Bold grid lines indicate which years’ policies were in the corpus and were annotated. The exact parameter counts displayed in this figure appear in Tables 56 in Appendix H.
Company Year Aim Attribute Condition Consequence Modality Recipient Sender Subject Total
Facebook 2005 1 7 5 0 1 6 4 5 29
Facebook 2006 31 42 48 11 32 40 28 30 262
Facebook 2007 31 50 56 15 39 50 33 31 305
Facebook 2008 34 54 58 18 39 56 38 33 330
Facebook 2009 62 78 82 31 51 70 52 41 467
Facebook 2010 63 89 102 40 61 79 57 51 542
Facebook 2011 66 87 100 39 64 80 61 54 551
Facebook 2015 26 55 42 7 24 44 41 26 265
Facebook 2016 29 55 39 8 24 44 43 28 270
Facebook 2017 30 56 41 7 24 44 41 27 270
Facebook 2019 51 94 64 17 49 77 70 62 484
NY Times 2001 17 22 22 8 8 21 15 16 129
NY Times 2002 18 20 21 5 7 21 16 14 122
NY Times 2003 27 27 24 8 11 28 16 15 156
NY Times 2004 31 30 34 9 14 30 18 17 183
NY Times 2005 34 34 37 9 16 35 22 23 210
NY Times 2006 37 36 36 11 17 37 25 22 221
NY Times 2007 9 10 10 1 7 11 6 6 60
NY Times 2008 7 10 9 1 9 10 5 6 57
NY Times 2009 7 10 11 1 7 11 6 6 59
NY Times 2010 7 11 9 1 7 11 6 6 58
NY Times 2011 8 14 11 2 10 14 7 9 75
NY Times 2012 58 76 69 22 36 69 46 41 417
NY Times 2013 55 66 64 22 37 59 39 37 379
NY Times 2014 53 64 63 25 39 62 40 35 381
NY Times 2015 44 52 50 17 34 47 30 29 303
NY Times 2016 43 54 50 18 35 47 30 27 304
NY Times 2017 45 58 55 20 36 54 31 30 329
NY Times 2018 44 52 57 21 25 49 40 35 323
NY Times 2019 48 57 59 20 25 53 41 37 340
GitHub 2008 6 8 8 3 4 7 8 4 48
GitHub 2009 5 8 9 4 4 8 8 4 50
GitHub 2010 5 8 7 3 5 8 8 4 48
GitHub 2012 4 8 8 2 5 7 6 3 43
GitHub 2013 6 8 9 4 5 9 7 3 51
GitHub 2014 5 7 9 4 5 8 6 3 47
GitHub 2016 5 9 9 3 5 8 8 4 51
GitHub 2019 52 91 90 26 57 81 45 51 493
Buzzfeed 2007 3 8 8 6 6 5 3 2 41
Buzzfeed 2008 3 7 8 4 6 5 3 2 38
Buzzfeed 2009 4 8 8 4 6 5 4 2 41
Buzzfeed 2011 4 7 8 4 6 5 4 2 40
Buzzfeed 2014 34 45 38 17 36 50 26 18 264
Buzzfeed 2015 33 44 40 18 36 48 25 19 263
Buzzfeed 2016 32 45 37 19 37 47 26 18 261
Buzzfeed 2019 59 71 49 20 38 54 29 41 361
Google 2010 1 0 1 0 0 0 1 0 3
Google 2011 1 0 0 0 0 0 0 0 1
Google 2015 35 44 44 19 27 37 18 22 246
Google 2016 36 45 45 16 27 34 19 24 246
Google 2017 38 47 44 18 28 36 19 24 254
Table 5. Counts of annotated parameters in the privacy policies of 10 prominent companies and organizations over time (continued on next page).
Company Year Aim Attribute Condition Consequence Modality Recipient Sender Subject Total
Bank of America 2005 6 17 9 2 12 9 10 8 73
Bank of America 2006 8 18 10 1 12 10 9 8 76
Bank of America 2011 30 26 33 16 20 31 21 11 188
Bank of America 2012 22 27 29 16 15 25 16 9 159
Bank of America 2014 29 29 28 14 24 27 23 9 183
Bank of America 2015 31 30 29 16 27 27 23 10 193
Bank of America 2016 31 29 27 16 27 26 22 8 186
Bank of America 2017 29 30 28 13 26 26 22 9 183
Bank of America 2018 31 29 27 14 26 27 23 8 185
Bank of America 2019 30 29 28 16 26 27 21 7 184
EFF 2000 4 10 14 2 8 10 7 3 58
EFF 2001 4 11 13 3 10 10 7 3 61
EFF 2002 5 7 11 3 5 7 3 7 48
EFF 2003 5 8 10 3 5 7 3 6 47
EFF 2004 4 5 12 4 4 7 4 7 47
EFF 2005 4 6 11 2 4 8 4 6 45
EFF 2006 15 32 32 10 21 26 12 15 163
EFF 2007 15 31 31 12 20 27 11 14 161
EFF 2008 16 29 32 11 20 26 9 16 159
EFF 2009 16 29 30 10 17 26 11 13 152
EFF 2010 16 27 31 12 18 23 13 15 155
EFF 2011 18 34 36 13 21 29 16 16 183
EFF 2012 17 34 35 6 22 31 16 16 177
EFF 2013 18 32 38 8 19 27 12 12 166
EFF 2014 20 34 34 11 21 30 14 14 178
EFF 2015 19 32 35 12 23 28 14 14 177
EFF 2016 22 33 34 13 24 28 15 14 183
EFF 2017 22 32 33 13 21 29 15 15 180
EFF 2018 21 32 34 14 21 31 13 14 180
EFF 2019 18 38 42 8 22 27 16 13 184
NSF 1999 4 10 7 1 1 6 3 2 34
NSF 2000 4 9 7 1 1 6 3 2 33
NSF 2001 4 11 8 1 1 6 2 3 36
NSF 2002 4 10 8 2 1 6 3 2 36
NSF 2003 4 10 8 1 1 6 3 2 35
NSF 2004 4 10 8 1 1 6 3 1 34
NSF 2005 3 10 8 1 1 6 3 2 34
NSF 2006 3 12 11 2 2 7 4 1 42
NSF 2007 3 10 11 2 3 7 4 2 42
NSF 2008 4 10 10 2 2 6 4 1 39
NSF 2009 4 11 10 1 1 6 3 1 37
NSF 2010 4 10 11 3 3 6 3 1 41
NSF 2011 11 17 21 10 6 11 6 9 91
NSF 2012 12 17 20 9 6 11 6 9 90
NSF 2013 11 16 21 7 6 10 7 9 87
NSF 2014 11 18 21 9 7 11 7 9 93
NSF 2015 12 17 21 8 6 11 6 9 90
NSF 2016 13 17 19 7 6 11 6 9 88
NSF 2017 11 20 21 9 6 12 8 12 99
NSF 2018 18 29 24 8 8 19 14 15 135
NSF 2019 18 28 23 8 7 17 14 15 130
Table 6. Counts of annotated parameters in the privacy policies of 10 prominent companies and organizations over time (continued from previous page)
Appendix I Parameter Variance Data
Website % Aim % Attribute % Condition % Consequence % Modality % Recipient % Sender % Subject Variance
mozilla.org 10.0 4.0 38.0 10.0 10.0 6.0 16.0 6.0 119.7
atlassian.net 4.2 12.5 25.0 33.3 8.3 8.3 4.2 4.2 119.0
doi.org 23.8 23.8 14.3 4.8 9.5 23.8 0.0 0.0 109.7
mozilla.com 10.4 6.2 35.4 8.3 10.4 6.2 16.7 6.2 98.0
apache.org 5.9 23.5 29.4 5.9 5.9 11.8 11.8 5.9 83.4
rlcdn.com 21.4 17.9 21.4 0.0 3.6 17.9 14.3 3.6 76.5
doubleverify.com 25.2 23.6 10.2 1.6 7.9 15.0 11.8 4.7 71.1
adsafeprotected.com 11.1 14.3 25.4 4.8 9.5 19.0 11.1 4.8 49.3
bbc.co.uk 10.0 16.7 23.3 3.3 6.7 20.0 10.0 10.0 46.8
who.int 8.8 17.6 23.5 10.3 4.4 17.6 11.8 5.9 43.3
tds.net 9.3 18.5 22.2 1.9 7.4 13.0 11.1 16.7 42.9
mit.edu 16.2 14.7 25.0 5.9 10.3 14.7 5.9 7.4 42.6
opera.com 13.1 17.9 22.0 5.4 13.7 16.7 7.7 3.6 41.7
nginx.com 23.4 13.2 20.4 6.4 11.9 11.1 7.2 6.4 41.0
office.net 14.9 19.9 19.9 2.5 8.7 16.8 10.6 6.8 40.5
att.net 16.7 16.7 14.8 1.9 9.3 22.2 9.3 9.3 39.9
rubiconproject.com 7.9 18.4 15.8 5.3 10.5 23.7 7.9 10.5 39.1
force.com 12.1 18.2 21.2 3.0 15.2 9.1 6.1 15.2 37.9
ibm.com 14.2 17.2 22.4 3.7 9.7 16.4 8.2 8.2 37.0
bbc.com 9.7 16.1 22.6 6.5 6.5 19.4 9.7 9.7 37.0
nih.gov 13.3 15.7 22.3 8.7 6.0 18.7 9.3 6.0 36.4
launchdarkly.com 13.9 19.8 18.8 7.9 5.9 17.8 10.9 5.0 35.3
europa.eu 16.7 8.3 20.8 16.7 8.3 8.3 16.7 4.2 34.7
criteo.com 17.1 20.7 15.9 4.9 11.0 15.9 8.5 6.1 32.6
adsrvr.org 11.3 15.1 20.8 3.8 13.2 18.9 7.5 9.4 32.5
android.com 12.0 17.1 18.8 4.3 16.2 16.2 11.1 4.3 32.3
webex.com 17.6 15.5 20.9 4.3 11.8 13.9 8.0 8.0 30.8
ubuntu.com 19.8 15.4 19.1 6.2 10.5 14.8 7.4 6.8 30.5
pinterest.com 14.1 18.6 18.8 3.9 7.2 16.6 11.9 8.9 30.4
cisco.com 17.7 17.2 19.8 5.2 10.9 14.1 8.3 6.8 30.1
salesforce.com 12.1 12.1 21.2 6.1 18.2 9.1 6.1 15.2 30.0
msn.com 17.1 20.2 16.5 5.7 7.0 15.5 9.0 9.1 29.5
windows.net 16.6 19.8 16.7 5.7 7.1 16.2 9.3 8.6 29.0
www.gov.uk 13.0 15.6 20.8 7.8 5.2 18.2 9.1 10.4 28.9
outlook.com 17.3 19.9 16.4 5.6 6.9 15.5 9.2 9.3 28.8
office365.com 17.3 19.9 16.5 5.6 7.3 15.4 9.0 9.0 28.8
skype.com 17.0 19.8 16.7 5.7 7.2 15.7 9.0 8.9 28.5
snapchat.com 5.8 18.5 16.0 6.2 14.5 18.5 12.7 7.6 28.3
bit.ly 11.1 19.2 18.8 5.4 11.5 16.9 10.7 6.5 27.9
live.com 17.0 19.7 16.6 5.7 7.4 15.6 9.0 9.1 27.9
casalemedia.com 12.9 18.6 15.7 1.4 11.4 17.1 11.4 11.4 27.8
windows.com 16.7 19.5 16.8 5.9 7.1 15.9 9.3 8.8 27.8
grammarly.com 16.8 17.7 18.2 4.1 12.3 14.5 9.5 6.8 27.7
w3.org 14.3 17.5 19.0 14.3 7.9 14.3 9.5 3.2 27.7
office.com 17.0 19.6 16.4 5.6 7.4 15.7 9.5 8.9 27.5
sharepoint.com 17.0 19.7 16.4 5.7 7.2 15.6 9.2 9.2 27.5
harvard.edu 16.4 17.9 13.4 3.0 11.9 17.9 7.5 11.9 27.3
name-services.com 17.2 13.8 6.9 6.9 13.8 20.7 13.8 6.9 27.0
google.com 13.6 19.9 18.2 5.5 11.4 15.3 7.2 8.9 26.8
microsoftonline.com 16.5 19.0 16.9 6.2 7.5 16.4 8.8 8.7 26.7
sentry.io 15.8 15.2 18.0 4.3 11.5 17.7 10.9 6.5 26.0
hubspot.com 13.2 17.1 20.6 5.7 9.2 15.8 10.5 7.9 25.8
unity3d.com 17.4 18.5 14.7 4.9 13.7 15.6 8.0 7.3 25.7
kaspersky.com 13.9 17.6 20.1 6.6 8.9 15.8 9.8 7.3 25.5
dnsmadeeasy.com 11.1 18.1 15.3 2.8 11.1 18.1 9.7 13.9 25.4
wal-mart.com 13.8 16.7 17.7 3.2 9.2 17.0 13.1 9.2 25.1
zemanta.com 19.0 18.0 15.0 8.0 10.0 15.0 10.0 5.0 24.9
googlevideo.com 14.2 19.7 17.2 5.4 11.3 15.5 7.1 9.6 24.8
goo.gl 14.3 20.0 18.0 6.5 10.6 14.3 6.9 9.4 24.6
shopify.com 15.3 16.5 16.5 2.7 9.2 17.2 11.6 11.1 24.5
cnn.com 17.2 13.1 14.2 4.1 18.7 13.9 12.0 6.7 24.3
sharethrough.com 13.7 18.8 16.4 4.7 14.6 15.3 10.0 6.6 24.3
flickr.com 11.1 20.7 17.3 5.8 11.5 15.4 9.1 9.1 24.2
azurewebsites.net 15.3 18.2 12.5 4.0 14.2 17.6 10.2 8.0 24.0
youtu.be 15.0 19.4 17.8 6.5 10.9 14.2 7.3 8.9 23.4
digicert.com 15.8 16.7 16.3 3.8 7.2 15.8 13.9 10.5 23.3
zoom.us 18.3 18.7 14.3 5.1 9.2 15.0 9.3 10.1 23.3
sourceforge.net 16.3 14.4 19.9 5.8 14.4 12.3 10.2 6.6 23.3
mcafee.com 12.7 18.7 13.3 3.0 12.7 17.5 12.7 9.6 23.1
ebay.com 10.5 16.3 22.1 11.6 7.0 12.8 8.1 11.6 23.1
facebook.com 10.6 19.5 12.9 3.5 9.8 16.2 14.6 12.9 22.8
smartadserver.com 18.3 18.3 12.7 6.3 11.1 16.7 8.7 7.9 22.6
wikimedia.org 17.3 16.8 18.5 7.0 8.8 14.0 10.5 7.0 22.6
reddit.com 7.1 17.2 19.2 6.1 13.6 15.7 9.6 11.6 22.6
roblox.com 15.3 16.7 18.1 4.0 13.9 14.1 8.7 9.3 22.5
mzstatic.com 14.6 19.2 13.6 4.9 14.6 16.0 9.1 8.0 22.5
googletagmanager.com 13.9 19.3 18.0 6.6 11.5 14.3 7.4 9.0 22.4
google.co.uk 13.6 19.8 18.1 7.0 10.7 14.0 7.4 9.5 22.3
scorecardresearch.com 14.5 15.3 9.9 6.9 16.8 19.8 8.4 8.4 22.3
youtube.com 14.0 19.3 17.7 6.2 11.5 14.4 7.4 9.5 22.3
Table 7. Website privacy policies ranked by the variance of the percentages of individual parameter types out of all annotations (continued on next page).
Website % Aim % Attribute % Condition % Consequence % Modality % Recipient % Sender % Subject Variance
pubmatic.com 16.08 15.50 14.04 3.51 16.67 15.79 9.65 8.77 22.17
ampproject.org 18.23 18.42 14.36 5.34 8.84 14.92 9.58 10.31 22.17
googleapis.com 14.34 18.85 18.44 6.97 10.66 14.34 7.79 8.61 21.98
bing.com 16.67 17.65 15.20 4.90 9.31 16.67 10.78 8.82 21.83
3lift.com 10.64 16.60 18.30 6.38 14.47 16.17 11.49 5.96 21.78
b-cdn.net 17.81 16.44 15.07 2.74 12.33 13.70 10.96 10.96 21.68
applovin.com 12.57 17.28 16.23 4.71 16.23 15.71 8.90 8.38 21.65
azure.com 14.36 17.68 14.36 3.87 13.81 17.13 9.94 8.84 21.62
xiaomi.com 17.92 17.38 17.20 5.73 10.75 13.08 9.68 8.24 21.52
wikipedia.org 16.70 16.13 18.98 7.78 8.16 14.61 10.44 7.21 21.51
fastly.net 16.89 17.43 18.23 5.36 9.38 12.87 9.92 9.92 21.50
bidswitch.net 10.58 12.50 19.23 3.85 17.31 13.46 11.54 11.54 21.40
shipt.com 10.78 17.65 17.65 5.39 11.27 16.18 13.73 7.35 21.35
amazon.com 11.67 17.22 20.56 8.33 9.44 15.00 10.00 7.78 21.34
googleadservices.com 13.68 19.23 17.52 6.41 10.68 14.96 7.69 9.83 21.22
gstatic.com 12.45 18.67 18.26 7.05 10.79 15.77 7.47 9.54 21.22
reuters.com 15.79 18.42 16.27 5.74 12.68 14.59 9.09 7.42 21.09
netflix.com 16.52 17.86 16.07 5.36 11.61 15.18 9.38 8.04 20.90
sciencedirect.com 12.77 16.67 20.57 5.67 10.64 12.77 8.87 12.06 20.87
github.io 11.41 17.72 18.53 5.70 11.61 16.09 8.35 10.59 20.80
booking.com 15.41 18.24 15.25 4.25 11.64 15.72 10.38 9.12 20.60
doubleclick.net 13.62 18.72 17.45 5.96 11.06 15.32 8.09 9.79 20.59
github.com 11.48 17.83 18.03 5.53 11.07 16.60 9.02 10.45 20.54
ui.com 13.64 18.56 12.50 6.06 15.15 16.67 10.98 6.44 20.42
theguardian.com 16.75 16.75 16.75 7.39 13.79 14.29 6.90 7.39 20.38
google-analytics.com 14.80 18.80 17.60 6.80 11.20 14.00 7.60 9.20 20.33
hotjar.com 13.01 17.07 17.07 3.25 10.57 15.45 12.20 11.38 20.20
yahoo.com 10.77 15.38 16.92 3.08 15.38 14.62 10.00 13.85 20.08
comcast.net 15.03 17.34 15.03 5.20 11.56 16.76 12.14 6.94 19.95
appsflyer.com 16.27 13.25 18.67 6.02 10.24 15.66 12.65 7.23 19.67
gandi.net 12.06 15.00 18.24 7.35 16.76 15.00 8.53 7.06 19.38
t.co 14.94 18.31 16.49 5.84 8.70 15.45 9.61 10.65 19.27
apple.com 15.09 17.54 14.39 5.26 14.39 15.79 9.47 8.07 18.75
wordpress.org 12.75 17.45 19.46 8.05 10.74 14.09 10.07 7.38 18.65
registrar-servers.com 14.81 15.56 17.04 6.67 6.67 17.04 11.85 10.37 18.41
hp.com 11.43 15.92 14.69 4.90 15.51 17.14 12.65 7.76 18.37
tiktok.com 10.96 18.26 14.61 3.65 11.87 15.53 12.79 12.33 18.28
frontapp.com 16.30 14.13 17.93 8.15 10.33 16.85 7.61 8.70 18.23
amazonaws.com 11.11 16.16 17.68 5.56 11.11 17.17 9.09 12.12 17.98
tinyurl.com 15.38 15.38 13.46 5.77 15.38 15.38 13.46 5.77 17.96
netflix.net 15.62 16.52 16.52 6.25 12.05 15.62 10.27 7.14 17.82
go.com 8.84 17.01 17.35 5.78 11.90 16.33 12.93 9.86 17.78
dailymail.co.uk 15.64 17.82 14.55 5.27 9.82 16.18 10.36 10.36 17.77
intuit.com 15.03 15.03 16.26 3.68 13.80 15.34 10.12 10.74 17.60
creativecommons.org 10.04 16.87 16.87 5.22 15.66 14.46 11.65 9.24 17.49
amazon.co.uk 15.35 15.35 17.21 6.51 9.77 16.28 7.44 12.09 17.40
icloud.com 15.02 18.03 13.73 6.01 13.30 15.88 10.30 7.73 17.18
vimeo.com 8.46 16.05 16.70 5.42 14.32 16.27 10.20 12.58 17.07
spotify.com 14.41 17.94 17.06 5.00 10.88 12.06 10.00 12.65 17.05
soundcloud.com 11.15 16.07 20.33 7.54 9.84 13.77 9.84 11.48 16.75
taboola.com 13.53 17.16 11.55 4.29 14.19 16.50 12.87 9.90 16.69
epicgames.com 9.09 15.34 19.60 7.10 15.06 12.50 11.36 9.94 16.23
researchgate.net 16.13 17.42 16.13 8.06 9.35 14.19 11.61 7.10 16.14
nytimes.com 13.43 16.72 17.31 6.57 7.76 15.52 11.94 10.75 16.00
weebly.com 16.21 15.17 18.28 7.59 11.72 13.79 9.31 7.93 15.96
aol.com 7.69 15.38 15.38 11.54 11.54 19.23 11.54 7.69 15.85
twitch.tv 10.34 15.52 16.09 5.75 14.94 16.67 12.07 8.62 15.83
macromedia.com 12.74 16.88 18.15 7.32 12.74 14.65 8.92 8.60 15.81
imdb.com 12.50 14.13 16.85 7.07 10.33 17.93 13.59 7.61 15.78
issuu.com 14.81 16.40 15.87 4.76 13.23 14.81 9.52 10.58 15.75
adobe.com 12.06 17.14 18.41 7.62 12.70 14.29 9.21 8.57 15.72
washingtonpost.com 13.53 15.84 14.19 5.28 17.16 14.19 10.56 9.24 15.17
forbes.com 14.08 14.93 16.90 6.48 9.86 16.34 13.24 8.17 15.01
deviantart.com 10.68 16.83 17.80 7.77 12.30 15.86 9.39 9.39 14.75
paypal.com 12.75 16.17 16.06 4.37 13.11 14.76 11.81 10.98 14.37
linkedin.com 13.09 15.88 16.99 5.85 10.31 16.16 10.86 10.86 14.24
cdc.gov 13.25 16.06 18.47 6.83 12.05 13.25 11.65 8.43 14.17
badoo.com 17.24 14.73 16.61 9.72 9.40 15.05 9.09 8.15 14.11
slideshare.net 12.91 16.21 16.76 5.49 10.44 15.66 11.54 10.99 14.08
wp.com 12.36 16.22 16.60 5.41 10.81 15.83 11.20 11.58 13.93
samsung.com 13.85 15.82 12.53 5.49 14.73 16.70 10.11 10.77 13.32
instagram.com 12.99 16.14 15.75 7.09 14.57 14.96 11.02 7.48 13.01
wordpress.com 13.14 15.69 16.27 5.69 10.98 15.88 10.98 11.37 12.65
gravatar.com 12.77 15.91 16.50 5.70 11.00 15.52 11.39 11.20 12.62
espn.com 8.63 16.31 15.59 8.63 15.35 14.63 12.71 8.15 12.22
tumblr.com 12.21 16.54 16.79 8.40 12.72 15.27 9.41 8.65 11.96
dropbox.com 13.27 15.49 15.49 7.96 13.27 16.37 10.18 7.96 11.51
cloudflare.com 11.45 17.18 17.62 7.93 11.89 12.78 10.13 11.01 11.20
medium.com 10.92 15.28 16.59 9.17 11.79 16.59 8.30 11.35 10.62
newrelic.com 12.16 14.86 18.02 6.76 11.71 12.61 11.26 12.61 10.19
hipages.com.au 12.28 12.28 12.28 7.02 14.04 17.54 12.28 12.28 8.30
Table 8. Website privacy policies ranked by the variance of the percentages of individual parameter types out of all annotations (continued from previous page).
Appendix J Parameter Density Data
Website #$\frac{\#\,\,Aims}{\#\,\,Sentences}$$\frac{\#\,\,Attributes}{\#\,\,Sentences}$$\frac{\#\,\,Conditions}{\#\,\,Sentences}$$\frac{\#\,\,Consequences}{\#\,\,Sentences}$$\frac{\#\,\,Modalities}{\#\,\,Sentences}$$\frac{\#\,\,Recipients}{\#\,\,Sentences}$$\frac{\#\,\,Senders}{\#\,\,Sentences}$$\frac{\#\,\,Subjects}{\#\,\,Sentences}$ Total # parameters / # Sentences
b-cdn.net 0.0774 0.0714 0.0655 0.0119 0.0536 0.0595 0.0476 0.0476 0.4345
facebook.com 0.0389 0.0716 0.0473 0.0130 0.0358 0.0595 0.0534 0.0473 0.3666
bing.com 0.0599 0.0634 0.0546 0.0176 0.0335 0.0599 0.0387 0.0317 0.3592
slideshare.net 0.0459 0.0576 0.0596 0.0195 0.0371 0.0557 0.0410 0.0391 0.3555
linkedin.com 0.0459 0.0557 0.0596 0.0205 0.0361 0.0566 0.0381 0.0381 0.3506
zoom.us 0.0641 0.0654 0.0500 0.0179 0.0321 0.0526 0.0327 0.0353 0.3500
ampproject.org 0.0635 0.0641 0.0500 0.0186 0.0308 0.0519 0.0333 0.0359 0.3481
sciencedirect.com 0.0437 0.0570 0.0704 0.0194 0.0364 0.0437 0.0303 0.0413 0.3422
snapchat.com 0.0183 0.0585 0.0505 0.0195 0.0459 0.0585 0.0401 0.0241 0.3154
t.co 0.0467 0.0572 0.0515 0.0183 0.0272 0.0483 0.0300 0.0333 0.3125
tumblr.com 0.0375 0.0508 0.0516 0.0258 0.0391 0.0469 0.0289 0.0266 0.3070
icloud.com 0.0437 0.0525 0.0400 0.0175 0.0387 0.0462 0.0300 0.0225 0.2913
spotify.com 0.0417 0.0519 0.0493 0.0145 0.0315 0.0349 0.0289 0.0366 0.2891
issuu.com 0.0427 0.0473 0.0457 0.0137 0.0381 0.0427 0.0274 0.0305 0.2881
samsung.com 0.0398 0.0455 0.0360 0.0158 0.0423 0.0480 0.0290 0.0309 0.2872
mzstatic.com 0.0420 0.0550 0.0390 0.0140 0.0420 0.0460 0.0260 0.0230 0.2870
creativecommons.org 0.0287 0.0482 0.0482 0.0149 0.0447 0.0413 0.0333 0.0264 0.2856
apple.com 0.0430 0.0500 0.0410 0.0150 0.0410 0.0450 0.0270 0.0230 0.2850
forbes.com 0.0398 0.0422 0.0478 0.0183 0.0279 0.0462 0.0374 0.0231 0.2826
azure.com 0.0401 0.0494 0.0401 0.0108 0.0386 0.0478 0.0278 0.0247 0.2793
reddit.com 0.0197 0.0478 0.0534 0.0169 0.0379 0.0435 0.0267 0.0323 0.2781
wp.com 0.0339 0.0445 0.0456 0.0148 0.0297 0.0434 0.0307 0.0318 0.2744
azurewebsites.net 0.0417 0.0494 0.0340 0.0108 0.0386 0.0478 0.0278 0.0216 0.2716
wordpress.com 0.0355 0.0424 0.0440 0.0154 0.0297 0.0429 0.0297 0.0307 0.2701
epicgames.com 0.0245 0.0414 0.0529 0.0192 0.0406 0.0337 0.0307 0.0268 0.2699
gravatar.com 0.0344 0.0429 0.0445 0.0154 0.0297 0.0418 0.0307 0.0302 0.2696
vimeo.com 0.0228 0.0432 0.0450 0.0146 0.0386 0.0438 0.0275 0.0339 0.2693
instagram.com 0.0350 0.0434 0.0424 0.0191 0.0392 0.0403 0.0297 0.0201 0.2691
shopify.com 0.0410 0.0443 0.0443 0.0072 0.0247 0.0462 0.0312 0.0299 0.2689
goo.gl 0.0377 0.0528 0.0474 0.0172 0.0280 0.0377 0.0183 0.0248 0.2640
google-analytics.com 0.0389 0.0494 0.0462 0.0179 0.0294 0.0368 0.0200 0.0242 0.2626
ebay.com 0.0274 0.0427 0.0579 0.0305 0.0183 0.0335 0.0213 0.0305 0.2622
adobe.com 0.0315 0.0447 0.0480 0.0199 0.0331 0.0373 0.0240 0.0224 0.2608
tiktok.com 0.0286 0.0476 0.0381 0.0095 0.0310 0.0405 0.0333 0.0321 0.2607
comcast.net 0.0392 0.0452 0.0392 0.0136 0.0301 0.0437 0.0316 0.0181 0.2605
macromedia.com 0.0331 0.0439 0.0472 0.0190 0.0331 0.0381 0.0232 0.0224 0.2599
google.co.uk 0.0353 0.0513 0.0470 0.0182 0.0278 0.0363 0.0192 0.0246 0.2596
youtu.be 0.0389 0.0504 0.0462 0.0168 0.0284 0.0368 0.0189 0.0231 0.2595
googlevideo.com 0.0366 0.0506 0.0442 0.0140 0.0291 0.0399 0.0183 0.0248 0.2575
gstatic.com 0.0321 0.0481 0.0470 0.0182 0.0278 0.0406 0.0192 0.0246 0.2575
flickr.com 0.0285 0.0532 0.0446 0.0149 0.0297 0.0396 0.0235 0.0235 0.2574
cdc.gov 0.0341 0.0413 0.0475 0.0176 0.0310 0.0341 0.0300 0.0217 0.2572
dropbox.com 0.0341 0.0398 0.0398 0.0205 0.0341 0.0420 0.0261 0.0205 0.2568
googleapis.com 0.0368 0.0483 0.0473 0.0179 0.0273 0.0368 0.0200 0.0221 0.2563
googletagmanager.com 0.0357 0.0494 0.0462 0.0168 0.0294 0.0368 0.0189 0.0231 0.2563
doubleverify.com 0.0645 0.0605 0.0262 0.0040 0.0202 0.0383 0.0302 0.0121 0.2560
youtube.com 0.0357 0.0494 0.0452 0.0158 0.0294 0.0368 0.0189 0.0242 0.2553
google.com 0.0345 0.0506 0.0463 0.0140 0.0291 0.0388 0.0183 0.0226 0.2543
microsoftonline.com 0.0415 0.0479 0.0426 0.0156 0.0189 0.0412 0.0221 0.0218 0.2516
doubleclick.net 0.0342 0.0470 0.0438 0.0150 0.0278 0.0385 0.0203 0.0246 0.2511
windows.net 0.0415 0.0498 0.0420 0.0143 0.0177 0.0406 0.0234 0.0215 0.2508
github.io 0.0286 0.0444 0.0464 0.0143 0.0291 0.0403 0.0209 0.0265 0.2505
googleadservices.com 0.0342 0.0481 0.0438 0.0160 0.0267 0.0374 0.0192 0.0246 0.2500
windows.com 0.0417 0.0487 0.0420 0.0147 0.0177 0.0395 0.0231 0.0219 0.2494
dailymail.co.uk 0.0389 0.0444 0.0362 0.0131 0.0245 0.0403 0.0258 0.0258 0.2491
github.com 0.0286 0.0444 0.0449 0.0138 0.0276 0.0413 0.0224 0.0260 0.2490
wordpress.org 0.0317 0.0433 0.0483 0.0200 0.0267 0.0350 0.0250 0.0183 0.2483
imdb.com 0.0309 0.0349 0.0417 0.0175 0.0255 0.0444 0.0336 0.0188 0.2473
bit.ly 0.0275 0.0473 0.0464 0.0133 0.0284 0.0417 0.0265 0.0161 0.2472
pinterest.com 0.0346 0.0455 0.0462 0.0095 0.0177 0.0408 0.0292 0.0217 0.2452
medium.com 0.0265 0.0371 0.0403 0.0222 0.0286 0.0403 0.0201 0.0275 0.2426
soundcloud.com 0.0269 0.0388 0.0491 0.0182 0.0237 0.0332 0.0237 0.0277 0.2413
live.com 0.0409 0.0474 0.0399 0.0136 0.0177 0.0375 0.0216 0.0220 0.2407
office.com 0.0410 0.0471 0.0393 0.0134 0.0177 0.0378 0.0228 0.0214 0.2406
sharepoint.com 0.0409 0.0473 0.0394 0.0138 0.0173 0.0375 0.0221 0.0222 0.2406
office365.com 0.0415 0.0479 0.0397 0.0135 0.0176 0.0369 0.0215 0.0216 0.2401
deviantart.com 0.0256 0.0404 0.0427 0.0186 0.0295 0.0380 0.0225 0.0225 0.2399
harvard.edu 0.0393 0.0429 0.0321 0.0071 0.0286 0.0429 0.0179 0.0286 0.2393
outlook.com 0.0414 0.0475 0.0393 0.0133 0.0165 0.0370 0.0221 0.0222 0.2392
skype.com 0.0405 0.0472 0.0398 0.0137 0.0172 0.0376 0.0215 0.0213 0.2388
msn.com 0.0409 0.0481 0.0393 0.0135 0.0166 0.0370 0.0215 0.0217 0.2388
applovin.com 0.0297 0.0408 0.0384 0.0111 0.0384 0.0371 0.0210 0.0198 0.2364
twitch.tv 0.0242 0.0363 0.0376 0.0134 0.0349 0.0390 0.0282 0.0202 0.2339
badoo.com 0.0402 0.0344 0.0387 0.0227 0.0219 0.0351 0.0212 0.0190 0.2332
yahoo.com 0.0250 0.0357 0.0393 0.0071 0.0357 0.0339 0.0232 0.0321 0.2321
android.com 0.0278 0.0397 0.0437 0.0099 0.0377 0.0377 0.0258 0.0099 0.2321
fastly.net 0.0392 0.0404 0.0423 0.0124 0.0218 0.0299 0.0230 0.0230 0.2320
booking.com 0.0354 0.0419 0.0350 0.0098 0.0267 0.0361 0.0238 0.0210 0.2298
xiaomi.com 0.0411 0.0399 0.0395 0.0132 0.0247 0.0300 0.0222 0.0189 0.2294
researchgate.net 0.0370 0.0399 0.0370 0.0185 0.0214 0.0325 0.0266 0.0163 0.2293
Table 9. Website privacy policies ranked by the total ratio of the number of annotated parameters to the number of sentences in the policy (continued on next page).
Website #$\frac{\#\,\,Aims}{\#\,\,Sentences}$$\frac{\#\,\,Attributes}{\#\,\,Sentences}$$\frac{\#\,\,Conditions}{\#\,\,Sentences}$$\frac{\#\,\,Consequences}{\#\,\,Sentences}$$\frac{\#\,\,Modalities}{\#\,\,Sentences}$$\frac{\#\,\,Recipients}{\#\,\,Sentences}$$\frac{\#\,\,Senders}{\#\,\,Sentences}$$\frac{\#\,\,Subjects}{\#\,\,Sentences}$ Total # parameters / # Sentences
paypal.com 0.0291 0.0369 0.0366 0.0100 0.0299 0.0337 0.0269 0.0251 0.2282
name-services.com 0.0391 0.0312 0.0156 0.0156 0.0312 0.0469 0.0312 0.0156 0.2266
dnsmadeeasy.com 0.0250 0.0406 0.0344 0.0063 0.0250 0.0406 0.0219 0.0312 0.2250
tds.net 0.0208 0.0417 0.0500 0.0042 0.0167 0.0292 0.0250 0.0375 0.2250
www.gov.uk 0.0291 0.0349 0.0465 0.0174 0.0116 0.0407 0.0203 0.0233 0.2238
ibm.com 0.0317 0.0383 0.0500 0.0083 0.0217 0.0367 0.0183 0.0183 0.2233
digicert.com 0.0344 0.0365 0.0354 0.0083 0.0156 0.0344 0.0302 0.0229 0.2177
weebly.com 0.0352 0.0329 0.0397 0.0165 0.0254 0.0299 0.0202 0.0172 0.2171
hotjar.com 0.0282 0.0370 0.0370 0.0070 0.0229 0.0335 0.0264 0.0246 0.2165
amazon.com 0.0252 0.0373 0.0445 0.0180 0.0204 0.0325 0.0216 0.0168 0.2163
registrar-servers.com 0.0321 0.0337 0.0369 0.0144 0.0144 0.0369 0.0256 0.0224 0.2163
scorecardresearch.com 0.0312 0.0329 0.0214 0.0148 0.0362 0.0428 0.0181 0.0181 0.2155
cnn.com 0.0371 0.0282 0.0306 0.0089 0.0403 0.0298 0.0258 0.0145 0.2153
roblox.com 0.0328 0.0358 0.0388 0.0085 0.0299 0.0303 0.0188 0.0201 0.2150
3lift.com 0.0228 0.0356 0.0392 0.0137 0.0310 0.0347 0.0246 0.0128 0.2144
unity3d.com 0.0371 0.0394 0.0313 0.0104 0.0293 0.0334 0.0172 0.0155 0.2136
w3.org 0.0304 0.0372 0.0405 0.0304 0.0169 0.0304 0.0203 0.0068 0.2128
opera.com 0.0278 0.0379 0.0467 0.0114 0.0290 0.0354 0.0164 0.0076 0.2121
taboola.com 0.0286 0.0363 0.0244 0.0091 0.0300 0.0349 0.0272 0.0209 0.2116
wikipedia.org 0.0350 0.0338 0.0398 0.0163 0.0171 0.0307 0.0219 0.0151 0.2098
tinyurl.com 0.0323 0.0323 0.0282 0.0121 0.0323 0.0323 0.0282 0.0121 0.2097
pubmatic.com 0.0337 0.0325 0.0294 0.0074 0.0349 0.0331 0.0202 0.0184 0.2096
ui.com 0.0285 0.0388 0.0261 0.0127 0.0316 0.0348 0.0229 0.0134 0.2089
office.net 0.0309 0.0412 0.0412 0.0052 0.0180 0.0348 0.0219 0.0142 0.2075
adsafeprotected.com 0.0230 0.0296 0.0526 0.0099 0.0197 0.0395 0.0230 0.0099 0.2072
intuit.com 0.0311 0.0311 0.0336 0.0076 0.0286 0.0317 0.0209 0.0222 0.2069
amazonaws.com 0.0229 0.0333 0.0365 0.0115 0.0229 0.0354 0.0187 0.0250 0.2062
netflix.net 0.0322 0.0340 0.0340 0.0129 0.0248 0.0322 0.0211 0.0147 0.2059
netflix.com 0.0340 0.0368 0.0331 0.0110 0.0239 0.0312 0.0193 0.0165 0.2059
smartadserver.com 0.0373 0.0373 0.0260 0.0130 0.0227 0.0341 0.0179 0.0162 0.2045
gandi.net 0.0246 0.0306 0.0373 0.0150 0.0343 0.0306 0.0174 0.0144 0.2043
nih.gov 0.0272 0.0319 0.0455 0.0177 0.0122 0.0380 0.0190 0.0122 0.2038
washingtonpost.com 0.0276 0.0323 0.0289 0.0108 0.0349 0.0289 0.0215 0.0188 0.2036
wikimedia.org 0.0353 0.0341 0.0377 0.0143 0.0179 0.0286 0.0214 0.0143 0.2036
ubuntu.com 0.0400 0.0312 0.0387 0.0125 0.0213 0.0300 0.0150 0.0138 0.2025
go.com 0.0179 0.0343 0.0350 0.0117 0.0240 0.0330 0.0261 0.0199 0.2019
cloudflare.com 0.0230 0.0346 0.0355 0.0160 0.0239 0.0257 0.0204 0.0222 0.2012
hubspot.com 0.0264 0.0343 0.0414 0.0114 0.0185 0.0317 0.0211 0.0158 0.2007
kaspersky.com 0.0279 0.0352 0.0402 0.0132 0.0178 0.0317 0.0197 0.0147 0.2005
wal-mart.com 0.0275 0.0332 0.0353 0.0064 0.0184 0.0339 0.0261 0.0184 0.1992
theguardian.com 0.0324 0.0324 0.0324 0.0143 0.0267 0.0277 0.0134 0.0143 0.1937
who.int 0.0170 0.0341 0.0455 0.0199 0.0085 0.0341 0.0227 0.0114 0.1932
espn.com 0.0166 0.0314 0.0300 0.0166 0.0295 0.0281 0.0244 0.0157 0.1923
sentry.io 0.0304 0.0292 0.0345 0.0083 0.0220 0.0339 0.0208 0.0125 0.1917
casalemedia.com 0.0245 0.0353 0.0299 0.0027 0.0217 0.0326 0.0217 0.0217 0.1902
rubiconproject.com 0.0150 0.0350 0.0300 0.0100 0.0200 0.0450 0.0150 0.0200 0.1900
sharethrough.com 0.0260 0.0355 0.0311 0.0090 0.0276 0.0289 0.0189 0.0124 0.1895
shipt.com 0.0204 0.0333 0.0333 0.0102 0.0213 0.0306 0.0259 0.0139 0.1889
mit.edu 0.0306 0.0278 0.0472 0.0111 0.0194 0.0278 0.0111 0.0139 0.1889
doi.org 0.0446 0.0446 0.0268 0.0089 0.0179 0.0446 0.0000 0.0000 0.1875
newrelic.com 0.0227 0.0277 0.0336 0.0126 0.0218 0.0235 0.0210 0.0235 0.1862
appsflyer.com 0.0301 0.0246 0.0346 0.0112 0.0190 0.0290 0.0234 0.0134 0.1853
hp.com 0.0208 0.0290 0.0268 0.0089 0.0283 0.0312 0.0231 0.0141 0.1823
nytimes.com 0.0245 0.0304 0.0315 0.0120 0.0141 0.0283 0.0217 0.0196 0.1821
mcafee.com 0.0230 0.0340 0.0241 0.0055 0.0230 0.0318 0.0230 0.0175 0.1820
bidswitch.net 0.0191 0.0226 0.0347 0.0069 0.0312 0.0243 0.0208 0.0208 0.1806
reuters.com 0.0282 0.0328 0.0290 0.0102 0.0226 0.0260 0.0162 0.0132 0.1783
grammarly.com 0.0298 0.0315 0.0323 0.0073 0.0218 0.0258 0.0169 0.0121 0.1774
sourceforge.net 0.0286 0.0254 0.0351 0.0101 0.0254 0.0217 0.0180 0.0115 0.1757
frontapp.com 0.0286 0.0248 0.0315 0.0143 0.0181 0.0296 0.0134 0.0153 0.1756
zemanta.com 0.0330 0.0312 0.0260 0.0139 0.0174 0.0260 0.0174 0.0087 0.1736
criteo.com 0.0292 0.0354 0.0271 0.0083 0.0187 0.0271 0.0146 0.0104 0.1708
amazon.co.uk 0.0259 0.0259 0.0291 0.0110 0.0165 0.0275 0.0126 0.0204 0.1690
bbc.com 0.0163 0.0272 0.0380 0.0109 0.0109 0.0326 0.0163 0.0163 0.1685
force.com 0.0200 0.0300 0.0350 0.0050 0.0250 0.0150 0.0100 0.0250 0.1650
salesforce.com 0.0200 0.0200 0.0350 0.0100 0.0300 0.0150 0.0100 0.0250 0.1650
nginx.com 0.0382 0.0215 0.0333 0.0104 0.0194 0.0181 0.0118 0.0104 0.1632
bbc.co.uk 0.0163 0.0272 0.0380 0.0054 0.0109 0.0326 0.0163 0.0163 0.1630
cisco.com 0.0281 0.0273 0.0315 0.0083 0.0174 0.0224 0.0132 0.0108 0.1589
webex.com 0.0273 0.0240 0.0323 0.0066 0.0182 0.0215 0.0124 0.0124 0.1548
mozilla.org 0.0142 0.0057 0.0540 0.0142 0.0142 0.0085 0.0227 0.0085 0.1420
att.net 0.0234 0.0234 0.0208 0.0026 0.0130 0.0312 0.0130 0.0130 0.1406
hipages.com.au 0.0172 0.0172 0.0172 0.0098 0.0196 0.0245 0.0172 0.0172 0.1397
mozilla.com 0.0142 0.0085 0.0483 0.0114 0.0142 0.0085 0.0227 0.0085 0.1364
launchdarkly.com 0.0179 0.0255 0.0242 0.0102 0.0077 0.0230 0.0140 0.0064 0.1288
europa.eu 0.0200 0.0100 0.0250 0.0200 0.0100 0.0100 0.0200 0.0050 0.1200
aol.com 0.0089 0.0179 0.0179 0.0134 0.0134 0.0223 0.0134 0.0089 0.1161
apache.org 0.0066 0.0263 0.0329 0.0066 0.0066 0.0132 0.0132 0.0066 0.1118
adsrvr.org 0.0112 0.0149 0.0205 0.0037 0.0131 0.0187 0.0075 0.0093 0.0989
rlcdn.com 0.0174 0.0145 0.0174 0.0000 0.0029 0.0145 0.0116 0.0029 0.0814
atlassian.net 0.0030 0.0089 0.0179 0.0238 0.0060 0.0060 0.0030 0.0030 0.0714
Table 10. Website privacy policies ranked by the total ratio of the number of annotated parameters to the number of sentences in the policy (continued from previous page).
Appendix K Visualizer Demonstration
Figure 18. Side-by-side comparison of Facebook’s 2015 and 2019 policies displayed using the visualizer, showing the GKC-CI annotations in the text. The legend for the color meanings is at the bottom of each GUI window. The comparison highlights how the information flow changes between the policies, with parameter annotations—indicated by text color—emphasizing the increase in specificity in information flows in the 2019 policy.